### v3

# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

Abstract
Detecting the significant features of the learning process of an artificial intelligence framework in the entire training and validation dataset can be determined as ’global’ explanations. Studies in the literature lack of accurate, low-complexity, and three-dimensional (3D) global explanations which are crucial in neuroimaging, a field with a complex representational space that demands more than basic two-dimensional interpretations. To fill this gap, we developed a novel explainable artificial intelligence (XAI) 3D-Framework that provides robust, faithful, and low-complexity global explanations. We evaluated our framework on various 3D deep learning networks trained, validated, and tested on a well-annotated cohort of 596 subjects from the TOP-OSLO study. The focus was on the presence and absence of the paracingulate sulcus, a variable feature of brain morphology correlated with psychotic conditions. Our proposed 3D-Framework outperforms traditional XAI methods in terms of faithfulness for global explanations. As a result, we were able to use these robust explanations to uncover new patterns that not only enhance the credibility and reliability of the training process but also reveal promising new biomarkers and significantly related sub-regions. For the first time, our developed 3D-Framework proposes a way for the scientific community to utilize global explanations to discover novel patterns in this specific neuroscientific application and beyond. This study can helps improve the trustworthiness of AI training processes and push the boundaries of our understanding by revealing new patterns in neuroscience and beyond.
Keywords XAI $\cdot$ Sulcus pattern  $\cdot$ PCS
1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. A recent study Mamalakis et al. (2024) has categorized the necessity for artificial intelligence (AI) explanations into four categories: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions and the stability of the evaluation protocol.
Specifically in neuroscience AI applications, where there exists significant variability in evaluation protocols and decision-making among experts, the inherent high uncertainty necessitates a greater need for explainability (non-explainable; Mamalakis et al. (2024)). To this end, applications in neuroscience usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, where AI has the potential to offer insights, and stabilize the validity and key aspects of the protocols Mitchell et al. (2023). This is particularly true for classification tasks where disease evaluation protocols are not yet firmly established (new-patterns discovery; Mamalakis et al. (2024)). For example, the folding of the human brain, established during gestation, is generally consistent among individuals. Post-natal development will affect sulcal metrics but the pattern of sulcation established at birth remains stable throughout life, preserving crucial prenatal neurodevelopmental information Cachia et al. (2016). These conditions makes them very interesting to study, but inter-individual variability create inherent variability in experts’ opinions and uncertainty in the evaluation protocol for tasks related to brain development and folding.
Detecting and annotating sulcal features from brain MRI is hindered by inter-individual variability in sulcation, and manual annotation is both time-consuming and subjective, impeding efficient use of extensive MRI databases. While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some secondary sulci present a more complex computational hurdle Borne et al. (2020). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental periods and may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues Jiang et al. (2021). For instance, the paracingulate sulcus (PCS), a variable secondary sulcus, is associated with cognitive performance and hallucinations in schizophrenia Fornito et al. (2006),Garrison et al. (2015),Gay et al. (2021),Ćurčić-Blake et al. (2023). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in structural MRI and frameworks that can provide both local and global explanations would allow for a more systematic assessment of its functional relevance and impact on critical functions in cognitive, behavioral, and developmental studies Simons et al. (2017).
However, an important challenge arises, particularly in neuroimaging, where the high representation dimensionality in AI applications, the high complexity, the involvement of unstable evaluation protocols, and the intricate expert annotations make these tasks challenging to explain and evaluate. To this end, ordinary 2D local or global XAI methods will fall short and may even be incorrect in some cases Mamalakis et al. (2024). Thus, there is an urgent need for 3D explanation frameworks that can surpass traditional XAI methods and deliver more reliable global explanations, which can be used to validate and identify new patterns in neuroscience.
Contribution. For the first time, we propose a comprehensive framework to validate and explain deep learning models in a way that reveals novel insights into pattern learning, thereby enhancing the credibility and reliability of these networks. To achieve this, we developed a novel XAI 3D-Framework, designed to assess the ability of deep learning networks to learn patterns in a binary classification task to detect the paracingulate sulcus, a highly variable secondary sulcus related to reality monitoring and psychotic conditions.
We applied two distinct local XAI methods, GradCam and SHAP, in 3D space to the entire dataset used in the binary classification task. A dimensionality reduction technique was then employed to extract global explanations from these XAI methods, as well as statistical patterns from the 3D brain inputs. By combining these two XAI methods with the statistical patterns in various configurations, we outperformed traditional XAI methods in global explainability, while also revealing the learning patterns within the network.
For the prediction of the paracingulate sulcus presence, a complex and high uncertainty computational hurdle, we developed two different 3D deep learning networks: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with variations in backbone options (2CNN-3D-MHL and simple-3D-MHL). The networks utilized 3D brain inputs derived from preprocessing structural MRI scans, which included grey-white surfaces and sulcal skeletons pre-processing images from both the right and left hemispheres. We used a well-annotated cohort of 596 subjects from the TOP-OSLO study Mørch-Johnsen et al. (2017) to train, validate, and test our deep learning networks, with a split of 70% for training, 20% for validation, and 10% for testing.
The proposed 3D framework surpasses traditional XAI methods in terms of faithfulness for global explanation. Importantly, our framework effectively identifies significant sub-regions by combining explainable features with statistical characteristics from the evaluation dataset. This approach offers a robust means to determine the patterns that contribute to a network’s decisions, thereby providing greater transparency, trust and further exploration in the outcomes of deep learning models used in neuroscience applications.
2 Related work
2.1 The application: Sulcal pattern studies
Cortical folding, which almost entirely develops in utero, results in an inter-individual variability that is often eschewed in populational studies. Yet, focusing on the variability of sulcal pattern can be motivated by different reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of mechanisms at play during neurodevelopment, and search for anatomo-functional correlates. Links to brain function can also be addressed to either investigate healthy functional variability (e.g. cingulate folding pattern and functional connectivity Fedeli et al. (2022)), or relationships between cortical folding pattern and pathological outcomes (e.g. paracingulate folding and hallucinations in schizophrenia Rollins et al. (2020)). The study of cortical folding variability can be addressed through global methods, with whole brain or regional considerations of generic sulcal parameters (e.g. gyrification index or sulcal pits), but most refined investigations of sulcal patterns require the capability for a focus on specific sulci, hence the need for automated sulcal recognition methods. A number of different methods have been developed for automated sulcal labelling in a general setting (for review, see Mangin et al. (2015)). Yet, to the best of our knowledge, no method is currently implemented to automatically label the paracingulate sulcus in a 3 dimensional approach (a 2 dimensional approach for a specific MRI slice has been proposed Yang et al. (2019)). The fact that the paracingulate sulcus is generally ommited in whole-brain labelling is probably related to the way in which it is anatomically specified: it is not only defined by its location in the brain but also by its orientation (parallel to the cingulate sulcus). Hence, even newer sulcal labelling implementations do not identify it within the folding of the medial frontal cortex Borne et al. (2020). Nevertheless, investigating the whole-brain anatomical correlates of sulcal variability incurred by the presence of the paracingulate sulcus is potentially important to understand some of the common and severe psychotic symptoms. Automatic detection of the paracingulate sulcus is in itself useful for large dataset exploration, especially since it has previously been related to functional variability in reality monitoring Simons et al. (2017). Additionally and unprecedentedly, adding an explainability component allows us to interrogate the anatomical covariates of the paracingulate sulcus gaining a borader vision on the mechanisms at which it may have a role.
The automated labelling of sulci allow shape analysis and pattern investigation of specific sulci on large datasets. These analyses can be led using different methods, on which the following are a general overview. First, simple metrics can be extracted automatically using the BrainVISA software Cointepas et al. (2001), such as the length or mean depth Mangin et al. (2020). Similarily, the local gyrification index can be derived from these automatically extracted metrics, defined as the ratio between a sulcus’s surface area and the outer cortical area Cachia et al. (2008). Yet, these metrics provide limited information on sulcal pattern. Automated methods providing more advanced sulcal pattern characterization have been proposed, as, for example, the comparison of positioning of the superior temporal sulcus between two cohorts Plaze et al. (2011). Another way of investigating sulcal pattern is by automatically capturing shape features. This can be done by looking for a predefined shape feature, such as a characteristic bulk in a sulcus Coulon et al. (2015) or a sulcus’s depth profile Leroy et al. (2015). Alternatively, this can also be investigated by using machine learning to automatically extract shape features from either a sulcus or a group of sulci Sun et al. (2016); De Vareilles et al. (2023); de Vareilles et al. (2023); Meng et al. (2020), with no prior expectation on the shape feature to capture. In addition to offering methodologies for investigating sulcal shape variability, these studies highlight our limited current understanding of the implications of sulcal pattern variability.
2.2 Explainable methods
Recent studies have seen a significant surge in the exploration of XAI within medical image analysis and neuroimaging domains Samek et al. (2021); van der Velden et al. (2022); Quellec et al. (2021); Mamalakis et al. (2021). XAI methodologies are broadly categorized into interpretable and post-hoc approaches. Interpretable methods focus on models that possess inherent properties such as simulatability, decomposability, and transparency, often linked to linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms van der Velden (2023). On the other hand, post-hoc methods are typically used with AI techniques to reveal nonlinear mappings within complex datasets Samek et al. (2021); van der Velden (2023).
A widely used post-hoc technique is Local Interpretable Model-Agnostic Explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e., in the close neighborhood of the prediction of interest Singh et al. (2020). Post-hoc techniques include both model-specific approaches that address specific nonlinear behaviors and model-agnostic approaches that explore data complexity Samek et al. (2021); van der Velden (2023). In computer vision, model-agnostic methods such as LIME and perturbation techniques are widely used, while model-specific methods encompass feature relevance, condition-based explanations, and rule-based learning Bach et al. (2015); Rajani et al. (2019); van der Velden (2023).
In medical imaging, explainable methods often focus on attribution and perturbation techniques Tjoa and Guan (2020). Attribution techniques like LIME, LRP, GRAD-CAM, and SHAP, they identify important features for a given prediction by assigning relevance scores to the input features. Perturbation techniques assess the sensitivity of an AI prediction to specific input features by systematically altering sub-groups of the input data Tjoa and Guan (2020); Singh et al. (2020). GRAD-CAM is notably prevalent among explainable methods in medical imaging due to its ease of application and understanding, as well as its ability to map significant features in the imaging space using the activations of the last convolutional layers Singh et al. (2020).
By advancing these explainable methodologies, we can better interpret complex models, enhancing their transparency and trustworthiness, particularly in critical applications like medical imaging and neuroimaging. An important gap in the literature is the absence of three-dimensional representation frameworks that can explain complex models, such as those used in neuroscience, by providing faithful global explanations. Such frameworks could offer more accurate interpretations than established approaches, potentially improving AI transparency and uncovering new patterns in significant sub-regions involved in predictions.
3 Methods and Background
In this section, we describe the different 3D classifiers used to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we developed a novel XAI 3D-Framework for the non-explainable and new-patterns discovery tasks (Mamalakis et al. (2024)). Specifically, we utilized two distinct explanation methods, belonging to the post-hoc family, SHAP and GradCam, expanded in three-dimensional space. We concatenate the outcomes obtained from the two explainable methods with the results derived from a statistical feature extraction model. The statistical model was a transparent dimensionality reduction algorithm applied to the inputs data of the validation cohort. This whole approach was intended to mitigate any potential biases and enhance the robustness of the pattern learning discovery.
3.1 Architectural design of deep learning networks for the classification task
We used two different deep learning model for the binary classification task (Fig. 1b.,c.). The first network was a 3D Convolutional Neural Network (CNN) with five levels (see Fig. 1b.). Each level incorporated a CNN block with a 3D convolution layer, a 3D max-pooling layer, and a batch normalization layer. In the first three levels, the 3D convolution layer employed 64, 128, and 256 filters, respectively (as shown in Fig. 1b.). The last level connected to a multi-layer perceptron (MLP) for the final prediction (as illustrated in Fig. 1a.). The MLP comprised three distinct perceptrons and two dropout layers to estimate epistemic uncertainty.
a
b
c
Figure 1: The input modalities and the architecture of simple-3D-CNN and the simple-3D-MHL networks. a, Input modalities illustrated on a right hemisphere coronal slice. More analytically the raw MRI of a given subject, the corresponding grey-white surface, and the corresponding sulcal skeleton. b, The architecture of simple-3D-CNN network with explanation of 3D Convolution layer (3D Conv) and 3D Max Pooling layer on the left. c, The three dimension MHL model with two different backbone choices, the full simple-3D-CNN (simple-3D-MHL) and the two level simple-3D-CNN layer (2CNN-3D-MHL).
For the second network (Fig. 1c.), we used a combination of multi-head attention layers (MHL) to focus on the global diversity and variation of a backbone output. To reduced the biased choice of only one backbone selection, we opted for two distinct backbone networks: (i) a 3D Convolution layer block with two level of 64 and 128 filters a Global Average Pooling, and a perceptron of 32400 hidden layers (2CNN-3D-MHL), and (ii) the straightforward 3D CNN network outlined previously. (Fig. 1b., simple-3D-MHL). We used the multi-head attention mechanism described in Vaswani et al. (2017) and as we used a two-head attention of the same input, we called this approach "self-attention". The two heads ($head_{c}$) are given by:
$head_{c}=AT(QW^{Q}_{c},KW^{K}_{c},VW^{V}_{c})$ (1)
where $c$ is the backbone output. There are two outputs as we used the backbone twice, and each output is connected to a perceptron with $N$ hidden layers (where $N$ is equal to the product of the weight and height of the input image). The $AT$ is the attention layer and it computed by:
$AT(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (2)
where the input matrix are combination of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Queries are packed together into matrix $Q$. The keys and values are also packed together into matrices $K$ and $V$. The output of the MHL network is given from:
$MHL=Concat(head_{c},head_{c})$ (3)
where c defines by the $backbone_{output}$. The output of the MHL was passed again from the MLP presented above to make the final prediction.
3.2 Extending explainability methods in 3D space
In this study, we used two distinct explainable techniques: the widely-utilized sensitivity local explainability technique in medical imaging applications known as the GradCam method (Selvaraju et al. (2017)), and a robust attribution explainability technique called SHAP (Lundberg and Lee (2017)). Previously emphasized is the significance of the three-dimensional space in neuroscience applications. Consequently, there arises a necessity to extend these 2D XAI methods into three dimensions. In the pursuit of computing the class-discriminative localization map encompassing width $w$, height $h$, and depth $d$ within a specific 3D brain MR image corresponding to a class $c$ (PCS or noPCS), the computation involves determining the gradient of the score for class $c$, denoted as $y^{c}$, in relation to the $o^{th}$ feature activation map ($A^{o}$) of the final convolution layer in each deep network. To determine the importance weights ($\alpha^{c}_{o}$) for each $o$ feature activation map, global average pooling is employed over the width ($i$), height ($j$), and depth ($k$) of each feature.
$\alpha^{c}_{o}=\frac{1}{Z}\sum_{i}{{\sum_{j}\sum_{k}\frac{dy^{c}}{dA_{ijk}^{o}}}}$ (4)
where $Z$ the summation of $i$, $j$ and $k$. Moreover, we used a weighted combination of forward activation maps and a $ReLU$ to deliver the final GradCam activation map.
$GradCam=ReLU(\sum_{o}{\alpha^{c}_{o}A_{ijk}^{o}})$ (5)
SHAP (SHapley Additive exPlanations; Lundberg and Lee (2017)) computes the attribution of each pixel of an input image for a specific prediction of a computer vision task. The attribution explainability methods follow the definition of additive feature attribution mainly as a linear function of:
$g(f,x)=\phi_{0}+\sum_{i=1}^{M}\phi_{i}x_{i}$ (6)
where $f$ is the prediction network, $g(f,x)$ is the explanation model, $\phi_{i}$ is the importance of each feature attribution ($\phi_{i}\in\mathbb{R}$), and $M$ is the number of simplified input features (pixels). Shapley value estimation is one of the main mathematical formulations that the SHAP algorithm used to assign an importance value to each feature, representing the effect on the model prediction of including that feature (attribution). If we define a subset $S$ of the total feature space ($F$) of an input 3D image $I$ ($i=1...N$, where $N$ is the number of samples in the dataset), and $x_{i}$ is a 3D matrix of width $w$ and height $h$ and depth $d$ for the $i^{th}$ sample, and $x_{S}$ is the subset of chosen features in the 3D space, then:
$\phi_{i}=\sum_{S/(i)}\frac{|S|!(|F|-|S|-1)!}{|F|!}[f_{S(i)}(x_{S(i)})-f_{S}(x_{S})]$ (7)
Here, $f_{S(i)}$ is a model trained with the presented $x_{s}$ features, and $f_{S}$ is another model trained with the features withheld. For our study, we used Deep SHAP Lundberg and Lee (2017) to describe our deep learning network models. In this approach, we used a chain rule and linear approximation as described in Lundberg and Lee (2017).
3.3 The XAI 3D-Framework
The main purpose of the XAI 3D-Framework is to provide global explanations that can lead to new pattern discovery. To achieve this, we used two XAI methods, GradCam and SHAP in 3D space to extract local explanations from the deep learning classifiers (simple-3D-CNN, simple-3D-MHL; Fig. 2 b.). Concurrently, we assigned faithfulness and complexity scores to evaluate the quality of the local explanations, ensuring their effectiveness reaches an acceptable threshold (see Supplementary material Table 2). Following that, we used Principal Component Analysis (PCA; Pearson (1901)) to derive global explanations from the two XAI methods. Moreover, PCA dimensionality reduction was applied to both the sulcal skeleton and the grey/white surface 3D brain inputs to identify a variety of global feature importances (feature-shape; see Fig. 3). We experimented with a range of different PCA number of components, to find the most optimal configuration in terms of time efficiency and the proportion of cohort variance captured (more than 80.0%). The optimal solution was found with six components. Based on a preliminary research, we found that the global explanations derived from each of the two XAI methods (GradCam and SHAP), yield better results in terms of faithfulness and complexity when we average the six components of the PCA analysis using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs total-SHAP, and PCA0-GradCam vs total-GradCam). Consequently, to generate the final global explanation represented from the GradCam (total-GradCam) and SHAP (total-SHAP), we apply the weighted average to the six PCA components (see Fig. 3). The same approach is applied in the feature-shape domain to determine the global statistical feature importance of the sulcal skeleton and the grey/white surface in the 3D brain inputs (total-Shape; see Fig. 3). The weighted average formulation is defined as follows:
$G(X,W)=\frac{\sum{w_{i}x_{i}}}{\sum{w_{i}}}$ (8)
where W is the weight tensor and the X is the pixel images tensor.
a
b
Figure 2: Classification task determination and the transition from local to global 3D explaination. a, Illustration of the no paracingulate sulcus condition (no PCS), and the paracingulate sulcus condition (PCS, in green line) on two left hemispheres 3D white matter reconstructions obtained with BrainVISA. The cingulate sulcus is coloured in yellow and blue and the callosal sulcus is coloured in purple. b, The 3D explainable framework that provides both local and global interpretations and explanations of our deep learning 3D classification network’s results. The ratio of the faithfulness and complexity metrics are compute in that stage. In this example we include only the GradCam explainability method just for simplicity.
3.4 The global explanation of 3D-Framework
To compute the global explanation for the 3D framework, we first manually align and rescale the two global XAI explanations (total-SHAP, total-GradCam) to the global statistical features (total-Shape). Then, we use equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. We conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which represent a fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey/white surface inputs of the total TOP-OSLO cohort, considering both the right and left hemispheres.
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. This allowed us to identify the most relevant features for each classification class by examining the overlap between the results of the two models. Identifying these features provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey/white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern for determining the presence or absence of the paracingulate sulcus. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas.
4 Experiments
4.1 Cohort’s description and pre-processing image analysis
We used the structural MRI of 596 participants from the TOP-OSLO study Mørch-Johnsen et al. (2017) for a binary classification task. The participants encompassed individuals within the unaffected control spectrum (262), those within the schizophrenia spectrum (183), and those within the bipolar disorder spectrum (151). The participants were imaged using a 1.5 T Siemens Magnetom Sonata scanner (Siemens Medical Solutions, Erlangen, Germany). The modality used was T1-weighted structural MRI, and the input files consisted of DICOM images. Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in Mitchell et al. (2023), with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and it is illustrated in Fig. 2 a. Image analysis techniques were applied to all slices to reduce artefacts. We used noise filters such as binomial deconvolution, Landweber deconvolution Vonesch and Unser (2008), and curvature anisotropic diffusion image filters Perona and Malik (1990) to reduce noise in the images.
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the Global PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracts the total overlapping image. This total overlapping image is then registered on a sulcal probabilistic atlas to unveil the model’s pattern for determining the presence or absence of the paracingulate sulcus.
4.2 Processing of two distinct images inputs
We initially processed the brain structural MRIs using the BrainVISA software Cointepas et al. (2001) and extracted two different images for our input: the grey-white surface and the sulcal skeleton. These were extracted from the raw MRI from a succession of bias correction, histogram analysis, brain segmentation, hemisphere separation, dichotomization of the white matter from the union of grey matter and cerebrospinal fluid, and skeletonization of the result, as detailed in Rivière et al. (2002). Specifically, the grey/white boundary was obtained by minimizing a Markov field and the segmentation used homotopic deformations of the hemisphere bounding box, resulting in the grey-white surface, where voxels are dichotomised into either grey or white. The skeleton was then derived from this object by applying a homotopic erosion embedding a watershed alogrithm, that preserves the initial topology, resulting in the sulcal skeleton. These two modalities were then used to train and evaluate our networks as well as our explainability methods. Fig. 1a,b,c. show the structural MRI and the corresponding grey-white surface and sulcal skeleton outputs from BrainVISA.
4.3 Hyper-parameters initialization
After randomly shuffling the data, each dataset was split into training, validation, and testing sets containing $70\%$, $20\%$, and $10\%$ of the total number of images, respectively. Sparse categorical cross-entropy was used as the cost function and the loss function was optimized using the Adam algorithm Kingma and Ba (2014). The learning rate was exponentially decreased during the first 50 epochs and then fixed at $0.0001$ for the last 50 epochs. To train the networks, an early stopping criterion of 10 consecutive epochs was employed and a maximum of 100 epochs was used for all input modalities (sulcal skeleton and grey/white matter) for both the left (L) and right (R) hemispheres. Finally, we use data augmentation techniques including rotation (around the center of the image by a random angle from the range $[-15^{\circ},15^{\circ}]$), width shift (up to $20$ pixels), height shift (up to $20$ pixels), and Zero phase Component Analysis (ZCA; Bell and Sejnowski (1997))whitening (add noise in each image) to avoid overfitting.
Evaluation metrics for the explanation
A crucial aspect of this study lies in evaluating how accurate and comprehensive is a local or global explanation. To derive a useful explanation, two primary scores play a pivotal role: faithfulness and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations Yeh et al. (2019). For a deep neural network $f$, and input features $\boldsymbol{x}$, the feature importance scores (also known as "attribution scores") are derived in a way such that when we set particular input features $\boldsymbol{x_{s}}$ to a baseline value $\boldsymbol{x_{s}^{f}}$, the change in the network’s output should be proportional to the sum of attribution scores of the perturbed features $\boldsymbol{x_{s}}$. We quantify this by using the Pearson correlation between the sum of the attributions of $\boldsymbol{x_{s}}$ and the difference in the output when setting those features to a reference baseline Bhatt et al. (2020). Thus, we define the faithfulness of an explanation method $g$ as:
$M_{faith}(f,g;\boldsymbol{x})=corr_{S}(\sum_{i\in S}g(f,\boldsymbol{x})_{i},f(\boldsymbol{x})-f(\boldsymbol{x}[\boldsymbol{x_{s}}=\boldsymbol{x_{s}^{f}}]))$ (9)
where $S$ is a subset of indices ($S\subseteq[1,2,3...d]$), $\boldsymbol{x_{s}}$ is a sub-vector of an input $\boldsymbol{x}$ ($\boldsymbol{x}=\boldsymbol{x_{s}}\cup\boldsymbol{x_{f}}$ and $\boldsymbol{x_{f}}$ the unchanged features of $\boldsymbol{x}$ image). The total number of the $\boldsymbol{x_{s}}$ sub-vectors, which partition an image is $d$. We denote as $\boldsymbol{x_{s}}$ the changed features, and $\boldsymbol{x_{f}}$ the unchanged features of $\boldsymbol{x}$ image.
If for an image $\boldsymbol{x}$, explanation $g$ highlights all $d$ features, then it may be less comprehensible and more complex than needed (especially for large $d$). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible Bhatt et al. (2020). If $P_{g}$ is a valid probability distribution and the $P_{g}(i)$ is the fractional contribution of feature $x_{i}$ to the total magnitude of the attribution, then we define the complexity of the explanation $g$ for the network $f$ as:
$M_{compx}(f,g;\boldsymbol{x})=\sum_{i=1}^{d}P_{g}(i)(log(\frac{1}{P_{g}(i)}))$ (10)
where:
$P_{g}(i)=\frac{|g(f,\boldsymbol{x})_{i}|}{\sum_{j=1}^{d}|g(f,\boldsymbol{x})_{j}|}$ (11)
In order to evaluate these two explainability metrics, we used the software developed by Hedström et al. (2023). This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for explanation methods. We note that we used a zero baseline (’black’; $\boldsymbol{x_{s}^{f}}$= $\boldsymbol{0}$), and 70 random perturbations to calculate the Faithfulness score.
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized the software developed by Hedström et al. (2023). As a first step we manually align and rescale the two global XAI explanations (total-SHAP and total-GradCam) to the global statistical features (total-Shape). The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanations served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual inputs brains of the cohorts are aligned in the same template, and the most significant variability of each class is assigned in the total-Shape. Consequently, we anticipate the classifier to predict correctly whether a patient has a PCS or not, as observed.
4.4 Data availability
This study used the datasets of the TOP-OSLO Mørch-Johnsen et al. (2017) which can be obtained from University of Oslo upon request, subject to a data transfer agreement.
4.5 Code availability
The code developed in this study is written in the Python programming language using Keras/TensorFlow (Python) libraries. For training and testing of deep learning networks, we have used an NVIDIA cluster with $4$ GPUs and $64$ GB RAM memory. The code is publicly available in https://github.com/ece7048/3Dsulci.
5 Results
5.1 Classifiers performance for presence or absence of paracingulate sulcus
For brevity, we discuss in the main manuscript only the simple-3D-MHL results which outperformed the two-level CNN backbone network (2CNN-3D-MHL). The analytical tables and results for 2CNN-3D-MHL can be found in the supplementary material subsection 2.1 (’Classification results of the 2CNN-3D-MHL’ ; Table 1).
The performance of the simple-3D-MHL network in the left hemisphere was higher (around 73.00% in all testing metrics and 74.10% in all validation metrics) than that in the right hemisphere (around 58.00% in all testing metrics and 63.10% in the validation metrics). For the simple-3D-CNN, the performance of the network in the left hemisphere was higher (around 72.90% in all testing metrics and 74.00% in all validation metrics) than that in the right hemisphere (around 56.00% in all testing metrics and 63.00% in the validation metrics). The analytical Figures and outcomes for the simple-3D-MHL and simple-3D-CNN networks are presented in supplementary material subsection 2.2 (’Additional global explainability methods and different components PCA results’; Supplementary Extended DATA Fig. 1 a.,b.).
This discrepancy in performance between the left and right hemisphere was to be expected for two reasons. First, the paracingulate sulcus is more prominent in the left than in the right hemisphere Paus et al. (1996); Yücel et al. (2001), including in psychopathological situations such as schizophrenia Garrison et al. (2015), implying that the pattern recognition problem is better defined on the left hemisphere. Moreover, the left paracingulate sulcus shows more associations with regional cortical thickness and sulcal depth than the right one Fornito et al. (2006), implying more covariability of anatomical features contained in either of our input modalities with the presence of the paracingulate sulcus in the left hemisphere than in the right one.
5.2 Global explanations and their PCA component results
In this study, we extracted and evaluated the explainable results from both networks to avoid biased observations and to investigate whether there is a clear cause-and-effect relationship between the quality of explanation and prediction performance. The underlying assumption is that better explanations could lead to better prediction performance. We presented the results of the simple-MHL network in the main manuscript, as it showed slightly higher performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Extended Data Fig. 2 and 3).
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey/white surface inputs show significant focus on the medial surface of the frontal lobe, thalamus and cingulate gyrus for detecting the presence of the paracingulate sulcus (PCS). Conversely, the absence of paracingulate sulcus (noPCS) is primarily detected by focusing on the thalamus and the anterior frontal lobe. In the left hemisphere significant focusing occurs on the thalamus following by the anterior front lobe. In the right hemisphere the region of significant focus is the thalamus and the anterior front lobe, following by the cingulate gyrus in some cases (PCA-Shape) (refer to Fig. 4).
Concerning the sulcal skeleton inputs on the left and right sides, Fig. 5.a,b display the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network. The neural network distinctly emphasize key regions in both hemispheres. In the left hemisphere, there’s notable focus on superior temporal sulcus, insula, and inferior central and precentral sulci for the PCS class (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on superior temporal sulcus, collateral fissure, calcarine fissure, parieto-occipital fissure, and sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere highlights specific areas, including superior temporal sulcus, insula, and inferior central and precentral sulci (Fig. 5a.). In the right hemisphere, the focus shifts to superior temporal sulcus, inferior central sulcus, internal parietal sulcus, and anterior cingulate sulcus (Fig. 5b.).
To validate the pattern indicating the presence or absence of the paracingulate sulcus, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Extended DATA Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey/white surface inputs, though the primary regions of focus remained consistent. Based on preliminary analysis (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam), we found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to when we used the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
a Refer to caption
b Refer to caption
Figure 4: Simple-3D-MHL results on the left and right hemisphere of white-grey mater brain inputs. a,b, show the explainability results for the PCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam results (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views.
a Refer to caption
b Refer to caption
Figure 5: Simple-3D-MHL results on the left and right hemisphere of sulcal skeleton brain inputs. a-b, show the explainability results for the noPCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views.
5.3 Global explanations from the 3D-Framework and the pattern learning results using grey/white surface inputs
For the simple-3D-MHL on grey/white surface inputs (Fig. 6a.), the right hemisphere shows focus on the medial view, in the thalamus and cingulate gyrus both in the PCS and - to a lesser extent - in the noPCS conditions. The left hemisphere, shows different focuses based on the PCS and noPCS conditions. The PCS decision mostly relies on the inferior temporal lobe and anterior inferior frontal lobe in the lateral view, and on the thalamus, both the anterior and posterior parts of the cingulate gyrus, the anterior frontal lobe, and the temporal lobe in the medial view. The noPCS condition, on the other hand, shows predominant focus on the inferior prefrontal lobe, and the temporal lobe in the lateral view, and on the thalamus and cingulate gyrus in the medial view.
For the simple-3D-CNN on grey/white surface inputs (refer to Supplementary material Extended DATA Fig. 6 a.), in the right hemisphere, PCS condition, the highlights in the lateral view are in the inferior frontal lobe and inferior temporal lobe in the PCS condition, and those in the medial view are in the thalamus, cingulate gyrus, anterior occipital lobe, and posterior temporal lobe. Still on the right hemisphere, in the noPCS condition, the focus is on the medial view in the thalamus, the cingulate gyrus, the frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focus on the same regions, laterally in the middle frontal lobe, and inferior and superior parietal lobe, and medially in the thalamus and in the cingulate, frontal and parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focuses on the lateral view of the posterior temporal lobe.
For both networks, in right hemisphere, the focus is essentially on the medial view (except for the PCS condition in the right hemisphere), with the main contributions in the thalamus and cingulate gyrus, suggesting a rather local explainability of the paracingulate sulcus’s presence. Conversely, on the left hemisphere, the focus is much more globally distributed, with strong contributions stemming from the lateral view, suggesting that the developmental mechanisms leading to the presence of the paracingulate sulcus affect the brain at a larger scale Fornito et al. (2006).
Table 1: Performance metrics of global explanation scores of faithfulness, and complexity for the global state of the art explanation methods; total-SHAP, total-GradCam and the propose 3D-Framework of the simple-3D-MHL network.
Explainable evaluation score of used explainable methods
XAI metrics XAI method Left sulcal skeleton Left white/grey Right sulcal skeleton Right white/grey
PCS / noPCS PCS / noPCS PCS / noPCS PCS / noPCS
Faithfulness ($\uparrow$) total-SHAP 0.105 / 0.20 0.092 / 0.113 0.103 / 0.102 0.073 / 0.088
total-GradCAM 0.044 / 0.09 0.143 / 0.164 0.067 / 0.085 0.119 / 0.129
3D-Framework 0.223 / 0.274 0.207 / 0.222 0.188 / 0.195 0.192 / 0.214
Complexity ($\downarrow$) total-SHAP 14.593 / 14.586 14.572 / 14.544 14.571 / 14.572 14.583 / 14.582
total-GradCAM 14.473 / 14.555 14.373 / 14.403 14.586 / 14.582 14.596 / 14.547
3D-Framework 14.584 / 14.563 14.582 / 14.587 14.579 / 14.573 14.587 / 14.587
a Refer to caption
b Refer to caption
c Refer to caption
Figure 6: Explainable methods scores and the extraction of the total overlapping pattern learning explanation of the left and right hemisphere and different inputs with expert’s observation. a shows the total overlapping pattern learning results for the right and left hemisphere of the brain for grey/white surface images of the simple-3D-MHL network. b shows the total overlapping pattern learning results for the right and left hemisphere of the brain for sulcal skeleton inputs of the simple-3D-MHL network. c shows the pattern learning results from the overlapping of the two deep learning networks (simple-3D-MHL, simple-3D-CNN), on the total overlapping pattern learning results of the lateral and medial view, based on experts’ observation. For the grey/white surface input we used the acronyms, Is: inferior surface, As: anterior surface, Ps: posterior surface, T: thalamus, H: hypothalamus, Fl: Frontal lobe, Ol: occipital lobe, Tl: temporal lobe, Pl: parietal lobe, Cc: corpus callosum, CG: cingulate gyrus. For the skeleton sulcal input we used the acronyms, STS: superior temporal sulcus, CF: calcarine fissure, POF: parieto-occipital fissure, SPS: sub-parietal sulcus, ICS: inferior central sulcus, CS: central sulcus, NA: none.
5.4 Global explanation from the 3D-Framework and the pattern learning results using sulcal skeleton inputs
A consistent method to overlay the outputs of the sulcal skeleton similar to the process described for the grey/white surface outputs was applied. For the simple-3D-MHL on sulcal skeleton inputs (Fig. 6b.), in the right hemisphere, the PCS and noPCS show similar trends, with lateral focuses on the superior temporal sulcus and the inferior precentral sulcus, and medial focuses on the interior parietal sulcus, the sub-parietal sulcus and both the inferior and superior rostral sulci (dimmed in the PCS condition). In the left hemisphere, the PCS and noPCS conditions show very different highlights. In the PCS condition, the main lateral focuses are in the posterior part of the inferior temporal sulcus, the occipital sulcus and the inferior frontal sulcus, and the medial focuses are in the rostral sulci (superior and inferior) and in the parietal sulci (interior, sub-parietal and temporo-occipital fissure). In the noPCS condition, the main crontributions are in the superior temporal sulcus and precentral sulcus in the lateral view, and in the sub-parietal sulcus in the medial view.
For the simple-3D-CNN on sulcal skeleton inputs (see Supplementary material Extended DATA Fig. 6 b.) both the right and the left hemispheres show similar trends in the PCS and noPCS conditions. In the right hemisphere, the lateral view is negligible compared to the medial view, where the main focus is on the callosal and parieto-occipital fissures and on the insula. In the left hemisphere, the main focus is found laterally in the superior temporal sulcus and medially in the collateral fissure and in the sub-parietal sulcus.
We thereafter identified the common regions between the two networks’ outputs and summarized them within Fig. 6c. Specifically, for the accurate detection of PCS within the skeleton sulcal hemisphere inputs, the pivotal sub-regions encompass the superior temporal sulcus, inferior precentral sulcus, calcarine fissure, parieto-occipital fissure, and sub-parietal sulcus. Conversely, when PCS is absent, the critical sub-regions within the right hemisphere skeleton sulcal hemisphere inputs encompass the superior temporal sulcus, inferior central sulcus, and sub-parietal sulcus (Fig. 6c). Transitioning to the left hemisphere, the skeleton sulcal hemisphere inputs underscore the significance of the superior temporal sulcus and sub-parietal sulcus. When PCS is not present, the important left hemisphere sulcal skeleton inputs comprise the superior temporal sulcus, inferior and central sulcus (Fig. 6c).
5.5 Ablation study of our 3D explainability framework
An ablation study in the simple-3D-MHL model was conducted to evaluate various combinations of global explanations (total-GradCam and total-SHAP) and global feature importance (total-Shape) (see Table 2). The best results were achieved by assigning the highest weight (’8’, 0.85) to total-Shape. This aligns with the idea that the feature importance of the inputs plays a crucial role in the explanations. For both hemispheres in the sulcal skeleton, a weight of ’5’, 0.5, given to total-SHAP, produced the highest faithfulness scores. In the white/grey surface, total-GradCam with the same weight yielded the best results (a weight of ’5’, 0.5). There were no significant differences observed in the complexity scores among the different combinations. The same patterns were observed for the simple-3D-CNN (Supplementary material; Table 2).
Table 2: Performance metrics assessing the global explanation scores of faithfulness and complexity were computed across various weight combinations assigned to the global XAI methods (total-SHAP, total-GradCam) and global feature extraction (total-Shape) within the proposed 3D-Framework for evaluating the global explanations of the simple-3D-MHL network. These combinations, denoted as 851, 815, 185, 158, 518, and 581, represent fixed-order representations of the assigned weights in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations, with weight values of 0.85 represented as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For instance, 3D-Framework-851 refers to the proposed 3D framework with weight values of 0.85 in total-Shape, 0.5 in total-SHAP, and 0.1 in total-GradCam.
Ablation study of the proposed 3D-Framework
XAI metrics XAI method Left sulcal skeleton Left white/grey Right sulcal skeleton Right white/grey
PCS / noPCS PCS / noPCS PCS / noPCS PCS / noPCS
Faithfulness 3D-Framework-851 0.223 / 0.274 0.166 / 0.197 0.188 / 0.195 0.172 / 0.182
3D-Framework-815 0.204 / 0.233 0.207 / 0.222 0.165 / 0.174 0.192 / 0.214
3D-Framework-185 0.105 / 0.122 0.118 / 0.124 0.053 / 0.133 0.113 / 0.133
3D-Framework-158 0.074 / 0.095 0.143 / 0.173 0.122 / 0.076 0.155 / 0.162
3D-Framework-518 0.115 / 0.106 0.125 / 0.152 0.144 / 0.134 0.136 / 0.142
3D-Framework-581 0.126 / 0.138 0.087 / 0.102 0.135 / 0.142 0.103 / 0.112
Complexity 3D-Framework-851 14.584 / 14.563 14.582 / 14.582 14.579 / 14.573 14.595 / 14.592
3D-Framework-815 14.594 / 14.595 14.582 / 14.587 14.582 / 14.576 14.587 / 14.587
3D-Framework-185 14.595 / 14.584 14.585 / 14.582 14.572 / 14.595 14.584 / 14.587
3D-Framework-158 14.592 / 14.583 14.587 / 14.583 14.573 / 14.594 14.593 / 14.592
3D-Framework-518 14.583 / 14.585 14.592 / 14.584 14.573 / 14.576 14.594 / 14.594
3D-Framework-581 14.592 / 14.586 14.593 / 14.592 14.573 / 14.587 14.597 / 14.592
5.6 Evaluation of the Global explanation from the 3D-Framework and the XAI methods and the pattern learning results
To evaluate whether the global explanation from the 3D-Framework is superior to those provided by SHAP or GradCam, we scored the explanations with respect to faithfulness and complexity (see Table 1). For the simple-3D-MHL network, our proposed 3D framework outperformed total-GradCam and total-SHAP in terms of faithfulness score in the left hemisphere, with values exceeding 0.21, compared to the lower scores of less than 0.16 for total-GradCam and less than 0.11 for total-SHAP. In the right hemisphere, our proposed 3D framework again outperformed total-GradCam and total-SHAP, achieving faithfulness scores over 0.18, compared to scores of less than 0.13 for total-GradCam and less than 0.10 for total-SHAP. Regarding the complexity score, the 3D framework achieved the second-best result, with total-GradCam having the lowest score in the left hemisphere and total-SHAP in the right hemisphere.
We went one step further exploring the sulcal skeleton output of the simple-3D-MHL network as it slightly outperformed the simple-3D-CNN ((see Supplementary material Table. 2 ii.) in the classification task and deliver better global explanations faithfulness and complexity scores than the simple-3D-CNN. We applied an affine registration to the total overlapping results of sulcal skeleton inputs from each hemisphere onto a probabilistic atlas of sulci Perrot et al. (2011). Fig. 7 illustrates the distribution of the most relevant voxels for outcome decision within sulcal probabilistic areas according to different thresholds. The decisions on the right hemisphere were highly focused on specific sulci, with up to three sulci contributing to the 20% threshold, which we retained as the lower threshold (blue). Conversely, the left hemisphere predictions were based on broader considerations, with a number of sulci already contributing to the decision based on the 5% threshold, which we retained as the lower threshold for the left hemisphere (blue). For both conditions and both hemisphere, a specific focus is set on the superior temporal sulcus and its posterior branches, and a smaller but consistent contribution of the internal parietal and sub-parietal sulci. The noPCS condition on the left hemisphere additionally focuses on the precentral sulcus and the Sylvian fissure. Interestingly, no specific focus is oriented towards the internal frontal sulcus (S.F.int), the probabilistic region in which the paracingulate sulcus is located when present.
a Refer to caption
b Refer to caption
Figure 7: A representation depicting the presence and absence patterns of the paracingulate sulcus on the left hemispheres of the brain’s sulcal skeleton a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Extended DATA Fig. 5., and NA: undefined.
6 Discussion
Explainability is essential in medical imaging and neuroimaging fields. The study Mamalakis et al. (2024) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. To that end we apply their proposed guidance in the task of the binary classification task related to secondary brain sulcal region highly related with psycotic patient and for the first time, we develop a new-patterns discovery XAI 3D-Framework.
The XAI 3D-Framework was developed to provide robust and faithful global and local explanations for a deep learning network in a binary classification task related to brain sulcal topology. We focused on the classification task of determining the presence or absence of the paracingulate sulcus and employed two different 3D deep learning networks: a simple deep 3D convolutional neural network and a two-headed attention layer network. The global explanations generated by our framework surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a robust interpretation of the deep networks for the binary classification task.
Our carefully designed and developed XAI 3D-Framework not only offers local and global interpretations but also identifies significant sub-regions involved in the existence and absence of the PCS, a highly variable secondary sulcus region highly correlated with psycotic patient. We discovered significant associations between specific anatomical regions of the brain through grey-white surface and sulcal skeleton inputs, that were distinct in the presence or absence of PCS. These findings enhance our understanding of the neural mechanisms related to the developmental origins of the PCS and provide valuable insights for future research in the field of neurological and psychiatric disorders. The regions of focus, both stemming from the grey/white matter interface and sulcal skeleton inputs, call for a globalized reflection about the neurodevelopmental mechanisms leading to the formation of a paracingulate sulcus (or absence thereof).
The focus which is decentralized from the region of definition of the paracingulate sulcus, suggests that the most relevant information about its presence of the paracingulate sulcus is located elsewhere, which is counter intuitive. Yet, this result may relay developmental insights regarding the mechanisms leading to the potential formation of the paracingulate sulcus De Vareilles et al. (2023). When looking at the sulcal skeleton results, while the timing of apparition of the internal parietal and sub-parietal sulci are unreported, the superior temporal and prefrontal sulci are both reported to first appear around 26 weeks of gestational age in the fetus Garel et al. (2001); Habas et al. (2012), while the paracingulate sulcus (which would be part of the secondary cingulate folding) only starts to appear around 31 weeks of gestational age Garel et al. (2001). The relationship between those sulci could therefore indicate an earlier chronology of the events orchestrating the apparition of a paracingulate sulcus. This could hint towards a greater genetic constraint on this sulcus as earlier developing sulci have been reported to show more heritability Pizzagalli et al. (2020).
Incidentally, the relevance of the Sylvian fissure in the results may imply an intertwining of gyrification and opercularization in the mechanisms leading to the formation of a paracingulate sulcus. This could be a potential explanation for the commonly reported leftward predominance of the left paracingulate sulcus, as, while gyrification shows a rightward developmental advance, opercularization has been reported to show a leftward advance De Vareilles et al. (2023). Therefore, the concurrent intervention of these two phenomenons could justify the asymmetrical patterning of the developing paracingulate sulcus and potentially its resulting form; in that case, the presence of this sulcus could be influenced by even earlier developmental events as opercularization starts before gyration and is importantly genetically regulated Mallela et al. (2020)
Our study introduces an automated and impartial annotation process, enhancing the possibilities for comprehensive explorations of sulcal development in relationship to cognitive and behavioral development. In the realm of deep learning and neuroscience, our contribution is a pioneering 3D explainable framework validating pattern learning in computer vision classifications tasks. Applied to paracingulate detection, our frameworks allowed to explore anatomical covariability and incidentally delve into the chronology of events intertwining structural and functional consequences. The effectiveness of our framework remains dependent on the quality and diversity of accessible data, and its performance could differ in alternate contexts. Additionally, while the incorporation of explainability techniques contributes to interpretability, it might not encompass all underlying intricacies of network behavior. The selection of the two local methods necessitates further exploration of alternative approaches to ensure comprehensiveness. The enhancement of classification outcomes demands the incorporation of additional information beyond structural MRI data to achieve improved accuracy.
To further investigate and incorporate more intricate strategies, the combination of XAI techniques and reduce dimensionality methods can explore overlapping aspects. Furthermore, complex approaches like t-SNE for non-linear dimensionality reduction can be employed to delve deeper into the subject. Moving forward, we envision expanding the framework’s applicability to other neurological conditions and classification tasks (such as shizophrenia or bipolar disorder), employing external datasets like BeneMIN datasets Lisiecka et al. (2015), Deakin et al. (2018), and further refining the interpretability techniques. Our study paves the way for a more systematic exploration of sulcal variability through the lens of deep learning, which can help explore cognitive and functional variability as well as pathological changes.
7 Conclusion
In this study, we developed a novel XAI 3D-Framework to deliver both local and global explanations for new-patterns discovery task in the secondary sulcus region PCS. The global explanations derived from this framework outperformed established methods such as SHAP and GradCam in terms of faithfulness, achieving the quality needed to discover new global patterns. Additionally, the three-dimensional representation of the proposed framework allows for the robust identification of complex new patterns within the intricate dimensionality of neuroscience applications. To accomplish this, the 3D-Framework enables the identification of sub-regions that are critical to decision-making through a post-fusion transformation of global explanations and statistical features.
Significantly, our framework has revealed notable correlations between specific brain sub-regions and the presence or absence of the paracingulate sulcus, offering deeper insights into underlying developmental mechanisms. This approach opens the door to more robust and reliable explanations in medical imaging and other complex computer vision tasks.
This work advances the fields of both deep learning and neuroscience, enabling automated and impartial annotations while illuminating intricate neuroanatomical relationships across diverse subregions. Our XAI 3D-Framework offers unparalleled insights, thus setting the stage for comprehensive exploration within the ever-evolving domains of deep learning and neuroscience.
Acknowledgment
All research at the Department of Psychiatry in the University of Cambridge is supported by the NIHR Cambridge Biomedical Research Centre (NIHR203312) and the NIHR Applied Research Collaboration East of England. The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care. This study was supported by funding from the Medical Research Council, grant number: MR/W020025/1. We acknowledge the use of the facilities of the Research Computing Services (RCS) of University of Cambridge, UK. GKM consults for ieso digital health. All other authors declare that they have no competing interests.