### v4

# An explainable three dimensional framework to uncover learning patterns: A unified look in variable sulci recognition

Abstract
The significant features identified in a representative subset of the dataset during the learning process of an artificial intelligence model are referred to as a ’global’ explanation. Three-dimensional (3D) global explanations are crucial in neuroimaging where a complex representational space demands more than basic two-dimensional interpretations. Curently, studies in the literature lack accurate, low-complexity, and 3D global explanations in neuroimaging and beyond. To fill this gap, we develop a novel explainable artificial intelligence (XAI) 3D-Framework that provides robust, faithful, and low-complexity global explanations. We evaluated our framework on various 3D deep learning networks trained, validated, and tested on a well-annotated cohort of 596 MRI images. The focus of detection was on the presence or absence of the paracingulate sulcus, a highly variable feature of brain topology associated with symptoms of psychosis. Our proposed 3D-Framework outperformed traditional XAI methods in terms of faithfulness for global explanations. As a result, these explanations uncovered new patterns that not only enhance the credibility and reliability of the training process but also reveal the broader developmental landscape of the human cortex. Our XAI 3D-Framework proposes for the first time, a way to utilize global explanations to discover the context in which detection of specific features are embedded, opening our understanding of normative brain development and atypical trajectories that can lead to the emergence of mental illness.
Keywords XAI $\cdot$ Sulcus pattern  $\cdot$ PCS
1 Introduction
In both medical imaging and neuroscience, explainability holds paramount importance. Recently, the study by Mamalakis et al. (2024) introduced the necessity of explanations in artificial intelligence (AI) healthcare applications, categorizing them into four types: self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery. This categorization is based on the variability of expert opinions, the stability of the evaluation protocol, and the dimensionality of the problem.
In neuroscience AI applications, there commonly exists significant variability in evaluation protocols and decision-making among experts. The inherent high uncertainty creates a greater need for explainability (as the framework falls in the non-explainable category; Mamalakis et al. (2024)). To this end, applications usually need both "local" methods, which provide explanations for each AI prediction separately, and "global" methods, which derive explanations for the decision-making of the AI across the entire dataset. Such variability underscores the importance of thorough investigation and validation of AI model behavior. Even among experienced professionals, knowledge gaps can persist, and this is where AI has the potential to offer insights and stabilize the validity of key aspects of the evaluation protocols (Mitchell et al. (2023)). This is particularly true for classification tasks where the key features of a disease are not yet firmly established (matching the new-patterns discovery category; Mamalakis et al. (2024)).
The folding of the human cortical surface occurs during the perinatal period and remains constant for an individual for the rest of their life, much like a fingerprint, preserving early neurodevelopmental information (Cachia et al. (2016)). Broadly, human brains share many features of cortical folding (sulci), however strong inter-individual variability creates inherent divergence in experts’ opinions when it comes to labelling the more variable sulci, impeding the effort to automatically label sulci in the regions showing most variability.
While, to date, automated methods excel in the detection of most sulci, the variable shape and presence/absence of some sulci present a more complex computational hurdle (Borne et al. (2020)). Successful automation through generalized, unbiased annotation would greatly aid studies focusing on brain folding variations, which are proxies for a critical developmental period with information that may relate to cognitive, behavioral, and developmental outcomes, along with psychiatric and neurological disorders. Brain folding is linked to brain function, and specific folding patterns correlate with susceptibility to neurological issues (Jiang et al. (2021)). In particular, the paracingulate sulcus (PCS), a highly variable sulcus, is associated with cognitive performance and hallucinations in schizophrenia (Fornito et al. (2006), Garrison et al. (2015), Gay et al. (2021), Ćurčić-Blake et al. (2023)). To this end, developing networks that can identify the presence or absence of PCS (or multiple PCS elements) in brain magnetic resonance imaging (MRI) and creating frameworks that can provide both local and global explanations would allow for a more systematic characterization and a more comprehensive understanding of whole-brain correlates related to its presence. This, in turn, would allow for both more precise and holistic assessments of its functional relevance and impact on brain functions (Simons et al. (2017)).
However, an important challenge arises, particularly in neuroimaging, where the high representation dimensionality in AI applications, the high anatomical complexity, the involvement of unstable evaluation protocols, and the intricate yet uncertain expert annotations make these tasks challenging to explain and evaluate. Ordinary 2D local or global XAI methods fall short and may even be incorrect in some cases (Mamalakis et al. (2024)). Thus, there is a need for 3D explanation frameworks that can surpass traditional XAI methods and deliver more reliable global explanations.
Contribution. For the first time, we propose a comprehensive framework to validate and explain deep learning models in a way that reveals novel insights into pattern learning, thereby enhancing the credibility and reliability of these networks. To achieve this, we developed a novel XAI 3D-Framework designed to assess the ability of deep learning networks to learn patterns in a binary classification task to detect the presence or absence of the PCS, a highly variable sulcus related to reality monitoring and psychotic conditions.
We applied two distinct local XAI methods, GradCam and SHAP, in 3D space to the entire dataset used in the binary classification task. A dimensionality reduction technique was then employed to extract global explanations from these XAI methods, as well as statistical patterns from the 3D brain inputs. By combining these two XAI methods with the statistical patterns in various configurations we outperformed traditional XAI methods in global explainability while revealing the learning patterns within the network.
For the detection of the PCS presence we developed two different 3D deep learning networks: a simple 3D convolutional neural network (simple-3D-CNN) and a two-headed attention layer network with variations in backbone options (2CNN-3D-MHL and simple-3D-MHL). The networks utilized 3D brain inputs derived from preprocessing structural MRI scans, which included grey-white surface boundaries and sulcal skeletons from both the right and left hemispheres. We used a well-annotated cohort of 596 subjects from the TOP-OSLO study cohort (Mørch-Johnsen et al. (2017)) to train, validate, and test our deep learning networks, with a split of 70% for training, 20% for validation, and 10% for testing.
The proposed 3D framework surpassed traditional XAI methods in terms of faithfulness for global explanation. Importantly, our framework effectively identified significant sub-regions by combining explainable features with statistical characteristics from the evaluation dataset. This approach offers a robust means to determine the patterns that contribute to a network’s decisions, thereby providing greater transparency, trust and further exploration of the outcomes of deep learning models used in neuroscience applications.
2 Related work
2.1 The application: Sulcal pattern studies
Cortical folding, which develops in the perinatal period (i.e. in the last few months of gestation and the first few months after birth) results in an inter-individual variability that is often eschewed in populational studies. Interest in the variability of sulcal patterns can be motivated by different reasons: strict descriptive anatomy, refinement of inter-subject registration, investigation of mechanisms at play during neurodevelopment, and a search for anatomo-functional correlates. Links to brain function can also be addressed to investigate healthy functional variability (e.g. cingulate folding pattern and functional connectivity Fedeli et al. (2022)), or relationships between cortical folding pattern and pathological outcomes (e.g. paracingulate folding and hallucinations in schizophrenia Rollins et al. (2020)). The study of cortical folding variability can be addressed through global methods, with whole brain or regional considerations of generic sulcal parameters (e.g. gyrification index or sulcal pits), but most refined investigations of sulcal patterns require the capability for a focus on specific sulci, hence the need for automated sulcal recognition methods. A number of different methods have been developed for automated sulcal labelling in a general setting (for review, see Mangin et al. (2015)). Yet, to the best of our knowledge, no method is currently implemented to automatically label the PCS in a 3D approach (a 2D approach for a specific MRI slice has been proposed Yang et al. (2019)). The fact that the PCS is generally omitted in whole-brain labelling is related to the way in which it is anatomically specified; not only by its location in the brain, but also by its orientation (parallel to the cingulate sulcus). Hence, even newer sulcal labelling implementations do not identify the PCS within the folding of the medial frontal cortex (Borne et al. (2020)). Nevertheless, investigating the whole-brain anatomical correlates of sulcal variability incurred by the presence of the PCS is potentially important to understand some of the most common and severe symptoms of psychosis. Automatic detection of the PCS is in itself useful for large dataset exploration, especially since it has previously been related to functional variability in reality monitoring (Simons et al. (2017)). Additionally and unprecedentedly, adding an explainability component allows us to interrogate the anatomical covariates of the PCS gaining a broader vision of the functional mechanisms in which it may have a role.
2.2 Explainable methods
Recent studies have seen a significant surge in the exploration of XAI within medical image analysis and neuroimaging domains (Samek et al. (2021); van der Velden et al. (2022); Quellec et al. (2021); Mamalakis et al. (2021)). XAI methodologies are broadly categorized into interpretable and post-hoc approaches. Interpretable methods focus on models that possess inherent properties such as simulatability, decomposability, and transparency, often linked to linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms (van der Velden (2023)). On the other hand, post-hoc methods are typically used with AI techniques to reveal nonlinear mappings within complex datasets (Samek et al. (2021); van der Velden (2023)).
A widely used post-hoc technique is Local Interpretable Model-Agnostic Explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e. in the close neighborhood of the detected structure (Singh et al. (2020)). Post-hoc techniques include both model-specific approaches that address specific nonlinear behaviors and model-agnostic approaches that explore data complexity (Samek et al. (2021); van der Velden (2023)). In computer vision, model-agnostic methods such as LIME and perturbation techniques are widely used, while model-specific methods encompass feature relevance, condition-based explanations, and rule-based learning (Bach et al. (2015); Rajani et al. (2019); van der Velden (2023)).
In medical imaging, explainable methods often focus on attribution and perturbation techniques (Tjoa and Guan (2020)). Attribution techniques like LIME as well as Layer-wise Relevance Propagation (LRP), Gradient-weighted Class Activation Mapping (GradCAM), and Shapley Additive Explanations (SHAP) identify important features for a given prediction by assigning relevance scores to the input features. Perturbation techniques assess the sensitivity of an AI prediction to specific input features by systematically altering sub-groups of the input data (Tjoa and Guan (2020); Singh et al. (2020)). GradCAM is notably prevalent among explainable methods in medical imaging due to its ease of application and understanding, as well as its ability to map significant features in the imaging space using the activations of the last convolutional layers (Singh et al. (2020)).
By advancing these explainable methodologies we can better interpret complex models, enhancing their transparency and trustworthiness, particularly in applications like medical imaging, and neuroimaging. An important gap in the literature is the absence of three-dimensional representation frameworks that can explain complex models, such as those used in neuroscience, by providing faithful global explanations. Such frameworks could offer more accurate interpretations than established approaches, potentially improving AI transparency and uncovering new patterns in significant sub-regions involved in classification and prediction.
3 Methods and Background
In this section, we describe the different 3D classifiers used to address the binary classification problem of determining the presence or absence of PCS. Furthermore, we develop a novel XAI 3D-Framework for the non-explainable and new-patterns discovery tasks (Mamalakis et al. (2024)). We utilize two distinct explanation methods, belonging to the post-hoc family SHAP and GradCam, expanded in three-dimensional space. We concatenate the outcomes obtained from the two explainable methods with the results derived from a statistical feature extraction model. The statistical model is a transparent dimensionality reduction algorithm applied to the input data of the validation cohort. This whole approach is intended to mitigate any potential biases and enhance the robustness of the pattern learning discovery.
3.1 Architectural design of deep learning networks for the classification task
We used two different deep learning models for the binary classification task (Fig. 1b.,c.). The first network was a 3D Convolutional Neural Network (CNN) with five levels (see Fig. 1b.). Each level incorporated a CNN block with a 3D convolution layer, a 3D max-pooling layer, and a batch normalization layer. In the first three levels, the 3D convolution layer employed 64, 128, and 256 filters, respectively (as shown in Fig. 1b.). The last level connected to a multi-layer perceptron (MLP) for the final prediction (as illustrated in Fig. 1a.). The MLP comprised three distinct perceptrons and two dropout layers to estimate epistemic uncertainty.
a
b
c
Figure 1: The input modalities and the architecture of simple-3D-CNN and the simple-3D-MHL networks. a, Input modalities illustrated on a right hemisphere coronal slice. In more detail, the raw MRI of a given subject, the corresponding grey-white surface, and the corresponding sulcal skeleton. b, The architecture of simple-3D-CNN network with explanation of 3D Convolution layer (3D Conv) and 3D Max Pooling layer on the left. c, The three dimension MHL model with two different backbone choices, the full simple-3D-CNN (simple-3D-MHL) and the two level simple-3D-CNN layer (2CNN-3D-MHL).
For the second network (Fig. 1c.), we used a combination of multi-head attention layers (MHL) to focus on the global diversity and variation of a backbone output. To reduce the biased choice of only one backbone selection, we opted for two distinct backbone networks: (i) a 3D Convolution layer block with two levels of 64 and 128 filters a Global Average Pooling, and a perceptron of 32400 hidden layers (2CNN-3D-MHL), and (ii) the straightforward 3D CNN network outlined previously (Fig. 1b., simple-3D-MHL). We used the multi-head attention mechanism described in (Vaswani et al. (2017)) and as we used a two-head attention of the same input ("self-attention"). The two heads ($head_{c}$) are given by:
$head_{c}=AT(QW^{Q}_{c},KW^{K}_{c},VW^{V}_{c})$ (1)
where $c$ is the backbone output. There are two outputs corresponding to the two uses of the backbone, and each output is connected to a perceptron with $N$ hidden layers (where $N$ is equal to the product of the weight and height of the input image). The $AT$ is the attention layer and it computed by:
$AT(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (2)
where the input matrix are combinations of queries and keys of dimension $d_{k}$, and values of dimension $d_{v}$. Queries are packed together into a matrix $Q$. The keys and values are also packed together into matrices $K$ and $V$. The output of the MHL network is given by:
$MHL=Concat(head_{c},head_{c})$ (3)
where $c$ defines by the $backbone_{output}$. The output of the MHL was passed again from the MLP presented above to make the final prediction.
3.2 Extending explainability methods in 3D space
We used two distinct explainable techniques: the widely-utilized sensitivity local explainability technique in medical imaging applications known as the GradCam method ((Selvaraju et al. (2017)), and a robust attribution explainability technique called SHAP ((Lundberg and Lee (2017)).
The significance of the 3D space in neuroscience applications emphasises a necessity to extend 2D XAI methods into three dimensions. In the pursuit of computing the class-discriminative localization map encompassing width $w$, height $h$, and depth $d$ within a specific 3D brain MRI corresponding to a class $c$ (PCS or noPCS), the computation involves determining the gradient of the score for class $c$, denoted as $y^{c}$, in relation to the $n^{th}$ feature activation map ($A^{n}$) of the final convolution layer in each deep network. To determine the importance weights ($\alpha^{c}_{n}$) for each $n$ feature activation map, global average pooling is employed over the width ($i$), height ($j$), and depth ($k$) of each feature.
$\alpha^{c}_{n}=\frac{1}{Z}\sum_{i}{{\sum_{j}\sum_{k}\frac{dy^{c}}{dA_{ijk}^{n}}}}$ (4)
where $Z$ the summation of $i$, $j$ and $k$. Moreover, we used a weighted combination of forward activation maps and a $ReLU$ to deliver the final GradCam activation map.
$GradCam=ReLU(\sum_{n}{\alpha^{c}_{n}A_{ijk}^{n}})$ (5)
SHAP computes the attribution of each pixel of an input image for a specific prediction of a computer vision task. Attribution explainability methods follow the definition of additive feature attribution mainly as a linear function of:
$g(f,x)=\phi_{0}+\sum_{i=1}^{M}\phi_{i}x_{i}$ (6)
where $f$ is the prediction network, $g(f,x)$ is the explanation model, $\phi_{i}$ is the importance of each feature attribution ($\phi_{i}\in\mathbb{R}$), and $M$ is the number of simplified input features (pixels). Shapley value estimation is one of the main mathematical formulations that the SHAP algorithm uses to assign an importance value to each feature, representing the effect on the model prediction of including that feature (attribution). If we define a subset $S$ of the total feature space ($F$) of an input 3D image ($i=1...N$, where $N$ is the number of samples in the dataset), and $x_{i}$ is a 3D matrix of width $w$ and height $h$ and depth $d$ for the $i^{th}$ sample, and $x_{S}$ is the subset of chosen features in the 3D space, then:
$\phi_{i}=\sum_{S/(i)}\frac{|S|!(|F|-|S|-1)!}{|F|!}[f_{S(i)}(x_{S(i)})-f_{S}(x_{S})]$ (7)
Here, $f_{S(i)}$ is a model trained with the presented $x_{s}$ features, and $f_{S}$ is another model trained with the features withheld. For our study, we used Deep SHAP (Lundberg and Lee (2017) to describe our deep learning network models. This approach uses a chain rule and linear approximation as described in (Lundberg and Lee (2017)).
3.3 The XAI 3D-Framework
The main purpose of the XAI 3D-Framework is to provide global explanations that can lead to the discovery of new patterns. To achieve this, we used two XAI methods, GradCam and SHAP in 3D space to extract local explanations from the deep learning classifiers (simple-3D-CNN, simple-3D-MHL; Fig. 2 b.). Concurrently, we assigned faithfulness and complexity scores to evaluate the quality of the local explanations ensuring that their effectiveness reached an acceptable threshold (see Supplementary material Table 2). Following that, we used Principal Component Analysis (PCA; (Pearson (1901)) to derive global explanations from the two XAI methods. PCA dimensionality reduction was applied to both the sulcal skeleton and the grey-white surface 3D brain inputs (analytical description of the 3D inputs please advice sub-section 4.2) to identify a variety of global feature importance (feature-shape; see Fig. 3). We experimented with a range of different numbers of PCA components to find the optimal configuration in terms of time efficiency and the proportion of cohort variance captured ($>80.0\%$). The optimal solution was found with six components.
Based on preliminary research, we found that the global explanations derived from each of the two XAI methods (GradCam and SHAP) yielded better results in terms of faithfulness and complexity when we averaged the six components of the PCA analysis using a weighted tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001] (see Supplementary Material Table 2; PCA0-SHAP vs total-SHAP, and PCA0-GradCam vs total-GradCam). Consequently, to generate the final global explanation represented from the GradCam (total-GradCam) and SHAP (total-SHAP), we applied the weighted average to the six PCA components (see Fig. 3). The same approach was applied in the feature-shape domain to determine the global statistical feature importance of the sulcal skeleton and the grey-white surfaces in the 3D brain inputs (total-Shape; see Fig. 3). The weighted average formulation is defined as follows:
$G(X,W)=\frac{\sum{w_{i}x_{i}}}{\sum{w_{i}}}$ (8)
where W is the weight tensor and the X is the pixel images tensor.
a
b
Figure 2: Classification task determination and the transition from local to global 3D explaination. a, Illustration of the no PCS condition (no PCS), and the PCS condition (PCS, in green line) on two left hemisphere 3D white matter reconstructions obtained with BrainVISA. The cingulate sulcus is coloured in yellow and blue and the callosal sulcus is coloured in purple. b, The 3D explainable framework that provides both local and global interpretations and explanations of our deep learning 3D classification network’s results. The ratio of the faithfulness and complexity metrics were computed at that stage. In this example we include only the GradCam explainability method for simplicity.
3.4 The global explanation of 3D-Framework
To compute the global explanation for the 3D framework, we first manually aligned and rescaled the two global XAI explanations (total-SHAP, total-GradCam) to the global statistical features (total-Shape). Then, we used equation 8, incorporating a three-component weight tensor of $[0.85,0.5,0.1]$. We conducted an ablation study to identify the optimal combination of the weight tensor. The different cases we examined were denoted as 851, 815, 185, 158, 518, and 581, which are fixed-order representations of the methods in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations. The weights correspond to 0.85 as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For example, 3D-Framework-851 refers to the proposed 3D-Framework with weight values of 0.85 for total-Shape, 0.5 for total-SHAP, and 0.1 for total-GradCam. These components were derived from both the sulcal skeleton and the grey-white surface inputs of the total dataset, considering both the right and left hemispheres.
To minimize potential bias that might occur by relying on only one deep learning network, we applied this approach to both the simple-3D-CNN and simple-3D-MHL networks. Identifying the significant features of the networks explanations provides insights into the mechanisms driving the network’s decision-making process. For enhanced clarity, the key brain sub-regions of interest, corresponding to the sulcal skeleton and grey-white surface, are visually depicted in Fig. 1a. Finally, the global explanation from the 3D framework was registered on a sulcal probabilistic atlas to illustrate the model’s pattern associated with determining the presence or absence of the PCS. The registration process involved affine transformations, including translation, rotation, scaling, and geometry adjustments, to align with the probabilistic atlas.
4 Experiments
4.1 Cohort’s description and pre-processing image analysis
We used the structural MRI of 596 participants from the TOP-OSLO study (Mørch-Johnsen et al. (2017)) for a binary classification task. The participants encompassed individuals with a diagnosis on the schizophrenia spectrum (183), on the bipolar disorder spectrum (151), and unaffected control participants (262). T1-weighted images were acquired using a 1.5 T Siemens Magnetom Sonata scanner (Siemens Medical Solutions, Erlangen, Germany).
Two experts, A.A. and H.V., performed image annotations, categorizing them into two classes: ’no paracingulate sulcus’ (noPCS) and ’paracingulate sulcus’ (PCS). This annotation was based on the protocol described in (Mitchell et al. (2023)) with further details available in the supplementary materials under ’1.1 PCS classification of TOP-OSLO’ and is illustrated in Fig. 2a.
4.2 Processing of two distinct images inputs
We initially processed the brain structural MRIs using the BrainVISA software (Cointepas et al. (2001)) extracting two images as inputs for the classifier: the grey-white surface and the sulcal skeleton. These were extracted from the raw MRI with an established protocol consisting of bias correction, histogram analysis, brain segmentation, hemisphere separation, dichotomization of the white matter from the union of grey matter and cerebrospinal fluid, and skeletonization of the result, as detailed in (Rivière et al. (2002)). Specifically, the grey-white surface was obtained by minimizing a Markov field and the segmentation used homotopic deformations of the hemisphere bounding box, resulting in the grey-white surface, where voxels are dichotomised into either grey or white. The skeleton was then derived from this object by applying a homotopic erosion embedding a watershed alogrithm that preserves the initial topology resulting in the sulcal skeleton. These two modalities were then used to train and evaluate our networks as well as for our explainability methods. Fig. 1a,b,c. show the structural MRI and the corresponding grey-white surface and sulcal skeleton outputs from BrainVISA.
Figure 3: The proposed global 3D-Framework explanation. A weighted averaging (Weight tensor: [0.85, 0.7, 0.5, 0.3, 0.1, 0.001]) of six PCA components produces the Global PCA image for PCA-Shape, PCA-GradCam, and PCA-SHAP. Following this, a weighted averaging (Weight tensor: [0.85, 0.5, 0.1]) of the three Global PCA overlapping images extracted the total overlapping image. This total overlapping image was then registered on a sulcal probabilistic atlas to unveil the model’s pattern for determining the presence or absence of the PCS.
4.3 Hyper-parameter initialization
After randomly shuffling the data, each dataset was split into training, validation, and testing sets containing $70\%$, $20\%$, and $10\%$ of the total number of images, respectively. Sparse categorical cross-entropy was used as the cost function and the loss function were optimized using the Adam algorithm (Kingma and Ba (2014)). After manual hyper-parameter searching the best learning rate was a value of $0.0001$. We utilized a strategy of exponentially decreased during the first 50 epochs and then fixed at $0.0001$ for the last 50 epochs. To train the networks, an early stopping criterion of 10 consecutive epochs was employed and a maximum of 100 epochs was used for both input modalities (sulcal skeleton and grey-white surface) for both the left (L) and right (R) hemispheres. Finally, we use data augmentation techniques including rotation (around the center of the image by a random angle from the range $[-15^{\circ},15^{\circ}]$), width shift (up to $20$ pixels), height shift (up to $20$ pixels), and Zero phase Component Analysis (ZCA; Bell and Sejnowski (1997)) whitening (add noise in each image) to avoid overfitting.
Evaluation metrics for the explanation
A crucial aspect of this study lies in evaluating how accurate and comprehensive were the local and global explanations. To derive a useful explanation, two primary scores play a pivotal role: faithfulness and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations (Yeh et al. (2019)). For a deep neural network $f$, and input features $\boldsymbol{x}$, the feature importance scores (also known as "attribution scores") were derived in a way such that when we set particular input features $\boldsymbol{x_{s}}$ to a baseline value $\boldsymbol{x_{s}^{f}}$, the change in the network’s output was proportional to the sum of attribution scores of the perturbed features $\boldsymbol{x_{s}}$. We quantified this by using the Pearson correlation between the sum of the attributions of $\boldsymbol{x_{s}}$ and the difference in the output when setting those features to a reference baseline (Bhatt et al. (2020)). Thus, we define the faithfulness of an explanation method $g$ as:
$M_{faith}(f,g;\boldsymbol{x})=corr_{S}(\sum_{i\in S}g(f,\boldsymbol{x})_{i},f(\boldsymbol{x})-f(\boldsymbol{x}[\boldsymbol{x_{s}}=\boldsymbol{x_{s}^{f}}]))$ (9)
where $S$ is a subset of indices ($S\subseteq[1,2,3...d]$), $\boldsymbol{x_{s}}$ is a sub-vector of an input $\boldsymbol{x}$ ($\boldsymbol{x}=\boldsymbol{x_{s}}\cup\boldsymbol{x_{f}}$ and $\boldsymbol{x_{f}}$ the unchanged features of $\boldsymbol{x}$ image). The total number of the $\boldsymbol{x_{s}}$ sub-vectors, which partition an image is $d$. We denote as $\boldsymbol{x_{s}}$ the changed features, and $\boldsymbol{x_{f}}$ the unchanged features of $\boldsymbol{x}$ image.
If for an image $\boldsymbol{x}$, explanation $g$ highlights all $d$ features, then it may be less comprehensible and more complex than needed (especially if $d$ is large). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible (Bhatt et al. (2020)). If $P_{g}$ is a valid probability distribution and the $P_{g}(i)$ is the fractional contribution of feature $x_{i}$ to the total magnitude of the attribution, then we define the complexity of the explanation $g$ for the network $f$ as:
$M_{compx}(f,g;\boldsymbol{x})=\sum_{i=1}^{d}P_{g}(i)(log(\frac{1}{P_{g}(i)}))$ (10)
where:
$P_{g}(i)=\frac{|g(f,\boldsymbol{x})_{i}|}{\sum_{j=1}^{d}|g(f,\boldsymbol{x})_{j}|}$ (11)
In order to evaluate these two explainability metrics, we used the software developed by (Hedström et al. (2023). This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for explanation methods. Note that we used a zero baseline (’black’; $\boldsymbol{x_{s}^{f}}$= $\boldsymbol{0}$) and 70 random perturbations to calculate the faithfulness score.
Finally, to extract the faithfulness and complexity scores of the global explanations for the total-SHAP, total-GradCam, and the proposed 3D-Framework, we utilized again the software developed by (Hedström et al. (2023)). As a first step we manually aligned and rescaled the two global XAI explanations (total-SHAP and total-GradCam) to the global statistical features (total-Shape). The input image consisted of the total-Shape results, while the total-GradCam, total-SHAP, and 3D-Framework global explanations served as reference explanations, respectively. In this context, the score of a global explanation makes sense, as the individual input brains of the cohort were aligned in the same template and the most significant variability of each class was assigned in the total-Shape. Consequently, we anticipated the classifier to classify correctly whether a MRI has a PCS or not.
4.4 Data availability
This study used the dataset of the TOP-OSLO cohort (Mørch-Johnsen et al. (2017)) which can be obtained from University of Oslo upon request, subject to a data transfer agreement.
4.5 Code availability
The code developed in this study is written in the Python programming language using pytorch, Keras, tensorflow (Python) libraries. For training and testing of deep learning networks, we have used an NVIDIA cluster with $4$ GPUs and $64$ GB RAM memory. The code is publicly available in https://github.com/ece7048/3Dsulci.
5 Results
5.1 Classifiers performance for presence or absence of paracingulate sulcus
For brevity, we discuss in the main manuscript only the simple-3D-MHL results which outperformed the two-level CNN backbone network (2CNN-3D-MHL). The analytical tables and results for 2CNN-3D-MHL can be found in the supplementary material subsection 2.1 (’Classification results of the 2CNN-3D-MHL’ ; Table 1).
The performance of the simple-3D-MHL network in the left hemisphere was higher (around 73.00% in all testing metrics and 74.10% in all validation metrics) than that in the right hemisphere (around 58.00% in all testing metrics and 63.10% in the validation metrics). For the simple-3D-CNN, the performance of the network in the left hemisphere was higher (around 72.90% in all testing metrics and 74.00% in all validation metrics) than that in the right hemisphere (around 56.00% in all testing metrics and 63.00% in the validation metrics). The analytical Figures and outcomes for the simple-3D-MHL and simple-3D-CNN networks are presented in supplementary material subsection 2.2 (’Additional global explainability methods and different components PCA results’; Supplementary Extended DATA Fig. 1 a.,b.).
The discrepancy in performance between the left and right hemispheres was to be expected for two reasons. First, the PCS is more prominent in the left than in the right hemisphere (Paus et al. (1996); Yücel et al. (2001)), including in psychopathological conditions such as schizophrenia (Garrison et al. (2015)). Furthermore, the left PCS has a greater number of associations with regional cortical thickness and sulcal depth than the right PCS (Fornito et al. (2006)), implying more covariability of anatomical features contained in either of our input modalities with the presence of the PCS in the left hemisphere than the right.
5.2 Global explanations and their PCA component results
We extracted and evaluated the explainability results from both networks to avoid biased observations and to investigate whether there was a clear cause-and-effect relationship between the quality of explanation and prediction performance. The hypothesis was that better prediction performance could lead to better explanation. We present the results of the simple-MHL network in the main manuscript as it had slightly better performance compared to the simple-3D-CNN. The results for the simple-3D-CNN are thoroughly detailed in the supplementary material subsection 2.2 (refer to ’Additional global explainability methods and different components PCA results’; Supplementary Extended Data Fig. 2 and 3).
The first component analysis of PCA explainability results of simple-3D-MHL networks on the left and right hemisphere of the grey-white surface inputs mainly focus on the frontal lobe (mostly inferior lateral and inferior medial), the cingulate gyrus, the temporal lobe, and occasionally the thalamus for detecting the presence or absence of the PCS. More specifically, for the detection of the presence of PCS (paracingulate sulcus; Fig. 4) in the left and right hemisphere the simple-MHL network focuses more in the frontal lobe (medial and inferior lateral), cingulate gyrus (mostly anterior), temporal lobe and sometimes thalamus. Contrary for the absence of PCS (No paracingulate sulcus; Fig. 4) in the left hemisphere the network focuses in the frontal lobe (mostly inferior medial and inferior frontal), the temporal lobe, the cingulate gyrus, occasionally the thalamus, and specifically for the PCA-GradCam, the corpus callosum. On the right hemisphere the simple-MHL network focuses in frontal lobe (medial and lateral), the temporal lobe, and the cingulate gyrus.
Fig. 5.a,b displays the comprehensive explanations of PCA-GradCam, PCA-SHAP, and PCA-Shape for the simple-3D-MHL network when using the sulcal skeleton inputs. The neural network shows distributed attention but still emphasizes some key regions in both hemispheres. In the left hemisphere, for the PCS class, there is focus on the superior temporal sulcus and its branches, the posterior sylvian fissure, and parts of the central and precentral sulci, with some attention on the cingulate sulcus and the medial frontal sulcus (containing the PCS) (Fig. 5a.). Conversely, in the right hemisphere for the same class, emphasis is on the superior and inferior temporal sulci, cingulate sulcus, medial frontal sulcus (containing the PCS), and the sub-parietal sulcus (Fig. 5b.). For the noPCS class, the left hemisphere shows similar yet less specific attention than for the PCS class, with an additional focus on the sylvian fissure and insula, and the anterior cingulate sulcus (Fig. 5a.). In the right hemisphere, additional focus is shifted to the anterior cingulate sulcus, the sub-parietal sulcus, and elements of the ventricules (Fig. 5b.).
To validate the pattern indicating the presence or absence of the PCS, we assessed the variability across all six components of the PCA. Notably, we observed differences in the six PCA components between the explanations and the global feature importance (refer to Supplementary Extended DATA Fig. 4). There were differences in the intensity and extent of regions highlighted between the sulcal skeleton and grey-white surface inputs, although the primary regions of focus remained consistent. Based on preliminary analysis (see Supplementary Material Table 2; PCA0-SHAP vs. total-SHAP, and PCA0-GradCam vs. total-GradCam), we found that the faithfulness and complexity score of the PCA’s first component for SHAP and GradCam methods performed poorly compared to the weighted average output of all six components as described in subsection 3.4 (total-SHAP, total-GradCam). Consequently, we used total-SHAP, total-GradCam, and total-Shape to compute the global explanation for the 3D framework.
a Refer to caption
b Refer to caption
Figure 4: Simple-3D-MHL results on the left and right hemisphere of grey-white surface brain inputs. a,b, show the explainability results for the PCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam results (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)).
a Refer to caption
b Refer to caption
Figure 5: Simple-3D-MHL results on the left and right hemisphere of sulcal skeleton brain inputs. a-b, show the explainability results for the noPCS class images of the first component among the six components of PCA for the total input modality (PCA-Shape), the total corresponding GradCam (PCA-GradCam), and the total corresponding SHAP results (PCA-SHAP). The feature’s importance (pixel attribution) varies from 0 (blue color) to 1 (red color), with high importance being 1 for the PCA-GradCam and PCA-Shape results. The orientation of the results are based on the medial anatomical views. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)).
5.3 Global explanations from the 3D-Framework and the pattern learning results using grey-white surface inputs
For the simple-3D-MHL on grey-white surface inputs (see Fig. 6a.c), in the right hemisphere the focus was on the frontal lobe, the insula, and parietal lobe in the PCS existence and, in the PCS absence (noPCS) on the temporal lobe, frontal lobe, cingulate gyrus and parietal lobe. In the left hemisphere, the PCS decision mostly relied on the cingulate gyrus, frontal lobe parietal lobe, and corpus collosum. The noPCS condition predominantly focused on the frontal lobe, and lateral temporal lobe and cingulate gyrus (see Fig. 6a.c).
For the simple-3D-CNN on grey-white surface inputs (refer to Supplementary material Extended DATA Fig. 6 a.) in the right hemisphere, highlighted the lateral inferior frontal lobe, and inferior temporal lobe in the PCS condition and the thalamus, cingulate gyrus, lateral anterior occipital lobe, and posterior temporal lobe. In the noPCS condition, the focus was on the thalamus, cingulate gyrus, the medial frontal lobe, and the posterior temporal lobe. In the left hemisphere, both conditions globally focused on the same regions: the lateral middle frontal lobe, and inferior and superior parietal lobe, and the thalamus and in the cingulate, frontal and medial parietal lobes, with a special focus on the anterior cingulate gyrus. The PCS condition additionally focused on the lateral view of the posterior temporal lobe.
For both networks, in the right hemisphere the focus was primarily on the medial aspect of the brain (except for the PCS condition in the right hemisphere) with the main contributions in the frontal lobe and cingulate gyrus, suggesting a rather constrained explainability of the PCS presence. Conversely, in the left hemisphere, the focus was much more broadly distributed, with strong contributions stemming from the lateral cortex, suggesting that the developmental mechanisms leading to the presence of the PCS are related to the wider development of the brain (Fornito et al. (2006)).
Table 1: Performance metrics of global explanation scores of faithfulness, and complexity for the global state of the art explanation methods; total-SHAP, total-GradCam and the propose 3D-Framework of the simple-3D-MHL network.
Explainable evaluation score of used explainable methods
XAI metrics XAI method Left sulcal skeleton Left white/grey Right sulcal skeleton Right white/grey
PCS / noPCS PCS / noPCS PCS / noPCS PCS / noPCS
Faithfulness ($\uparrow$) total-SHAP 0.105 / 0.20 0.092 / 0.113 0.103 / 0.102 0.073 / 0.088
total-GradCAM 0.044 / 0.09 0.143 / 0.164 0.067 / 0.085 0.119 / 0.129
3D-Framework 0.223 / 0.274 0.207 / 0.222 0.188 / 0.195 0.192 / 0.214
Complexity ($\downarrow$) total-SHAP 14.593 / 14.586 14.572 / 14.544 14.571 / 14.572 14.583 / 14.582
total-GradCAM 14.473 / 14.555 14.373 / 14.403 14.586 / 14.582 14.596 / 14.547
3D-Framework 14.584 / 14.563 14.582 / 14.587 14.579 / 14.573 14.587 / 14.587
a Refer to caption
b Refer to caption
c Refer to caption
Figure 6: Explainable method’s scores and the extraction of the total overlapping pattern learning explanation of the left and right hemispheres and different inputs with expert’s observation. a shows the total overlapping pattern learning results for the right and left hemisphere of the brain for grey-white surface images of the simple-3D-MHL network. b shows the total overlapping pattern learning results for the right and left hemispheres of the brain for sulcal skeleton inputs of the simple-3D-MHL network. All the presented results are align and mapping in the ICBM 2009a Nonlinear Asymmetric atlas (Fonov et al. (2011, 2009)). c shows the pattern learning results from the overlapping of the simple-3D-MHL deep learning networks on the total overlapping pattern learning results of the lateral and medial views based on experts’ observation. For the grey-white surface input we used the acronyms, T: thalamus, H: hypothalamus, Fl: Frontal lobe, Ol: occipital lobe, Tl: temporal lobe, Pl: parietal lobe, Cc: corpus callosum, Cg: cingulate gyrus, NA: none. For the skeleton sulcal input we used the acronyms, S.T.s: superior temporal sulcus, S.T.i: inferior temporal sulcus, F.C.M.ant: anterior cingulate sulcus, S.pa.int:internal parietal sulcus, F.C.L.p: sylvian fissure, S.F.sup: superior frontal sulcus, S.F.int: internal frontal sulcus, S.F.inf: inferior frontal sulcus, S.Call: callosal sulcus.
5.4 Global explanation from the 3D-Framework and the pattern learning results using sulcal skeleton inputs
A consistent method to overlay the outputs of the sulcal skeleton, similar to the process described for the grey-white surface outputs, was applied. For the simple-3D-MHL on sulcal skeleton inputs (Fig. 6b,c), in the right hemisphere, the presence of PCS focuses on the lateral superior temporal sulcus, sylvian fissure, inferior frontal sulcus, and the anterior cingulate sulcus. The absence of PCS (noPCS; Fig. 6c) focuses on the inferior temporal sulcus, superior temporal sulcus, ventricule, callosal sulcus, and inferior frontal sulcus. In the left hemisphere, the PCS and noPCS conditions had very different highlights. In the PCS condition, the main focuses were on the lateral posterior part of the lateral superior temporal sulcus, posterior inferior temporal sulcus, sylvian fissure, and the inferior parietal sulcus. In the noPCS condition, the main contributions were in the sylvian fissure, lateral superior temporal sulcus, superior frontal sulcus, internal parietal sulcus, anterior cingulate sulcus, and ventricle.
For the simple-3D-CNN on sulcal skeleton inputs, aimed at the accurate detection of PCS within the sulcal hemisphere inputs, the pivotal sub-regions encompassed the superior temporal sulcus, inferior precentral sulcus, sylvian fissure and sub-parietal sulcus (see Supplementary Material Extended DATA Fig. 6b). Conversely, when PCS was absent, the critical sub-regions within the right hemisphere sulcal skeleton inputs encompassed the ventricule, superior temporal sulcus, internal parietal sulcus and rostral sulcus. Transitioning to the left hemisphere, the sulcal skeleton inputs underscore the significance of the superior temporal sulcus the, sylvian fissure, and the internal parietal sulcus. When PCS was not present, the important left hemisphere sulcal skeleton inputs comprised the superior temporal sulcus, ventricule, inferior precentral sulcus, internal parietal sulcus (see Supplementary Material Extended DATA Fig. 6b).
We thereafter identified the common regions between the two networks’ outputs. The overlap results of the two networks, simple-3D-CNN and simple-3D-MHL, for the presence and absence of PCS (noPCS) reveal several common regions of interest in both the right and left hemispheres. For the presence of PCS in the right hemisphere, both networks highlight the lateral superior temporal sulcus and sylvian fissure. In the left hemisphere under the PCS condition, both models emphasize the superior temporal sulcus, the sylvian fissure, and the internal parietal sulcus. For the absence of PCS (noPCS) in the left and right hemisphere, the networks overlap in highlighting the ventricule, part of superior temporal sulcus and the internal parietal sulcus.
5.5 Ablation study of our 3D explainability framework
An ablation study in the simple-3D-MHL model was conducted to evaluate various combinations of global explanations (total-GradCam and total-SHAP) and global feature importance (total-Shape) (see Table 2). The best results were achieved by assigning the highest weight (0.85) to total-Shape. This aligns with the idea that the feature importance of the inputs plays a crucial role in the explanations. For both hemispheres in the sulcal skeleton, a weight of 0.5 given to total-SHAP produced the highest faithfulness scores. For the grey-white surface inputs, total-GradCam with the same weight yielded the best results (a weight of 0.5). There were no significant differences observed in the complexity scores among the different combinations. The same patterns were observed for the simple-3D-CNN (Supplementary material; Table 2).
Table 2: Performance metrics assessing the global explanation scores of faithfulness and complexity were computed across various weight combinations assigned to the global XAI methods (total-SHAP, total-GradCam) and global feature extraction (total-Shape) within the proposed 3D-Framework for evaluating the global explanations of the simple-3D-MHL network. These combinations, denoted as 851, 815, 185, 158, 518, and 581, represent fixed-order representations of the assigned weights in the following sequence: total-Shape, total-SHAP, and total-GradCam explanations, with weight values of 0.85 represented as ’8’, 0.5 as ’5’, and 0.1 as ’1’. For instance, 3D-Framework-851 refers to the proposed 3D framework with weight values of 0.85 in total-Shape, 0.5 in total-SHAP, and 0.1 in total-GradCam.
Ablation study of the proposed 3D-Framework
XAI metrics XAI method Left sulcal skeleton Left white/grey Right sulcal skeleton Right white/grey
PCS / noPCS PCS / noPCS PCS / noPCS PCS / noPCS
Faithfulness 3D-Framework-851 0.223 / 0.274 0.166 / 0.197 0.188 / 0.195 0.172 / 0.182
3D-Framework-815 0.204 / 0.233 0.207 / 0.222 0.165 / 0.174 0.192 / 0.214
3D-Framework-185 0.105 / 0.122 0.118 / 0.124 0.053 / 0.133 0.113 / 0.133
3D-Framework-158 0.074 / 0.095 0.143 / 0.173 0.122 / 0.076 0.155 / 0.162
3D-Framework-518 0.115 / 0.106 0.125 / 0.152 0.144 / 0.134 0.136 / 0.142
3D-Framework-581 0.126 / 0.138 0.087 / 0.102 0.135 / 0.142 0.103 / 0.112
Complexity 3D-Framework-851 14.584 / 14.563 14.582 / 14.582 14.579 / 14.573 14.595 / 14.592
3D-Framework-815 14.594 / 14.595 14.582 / 14.587 14.582 / 14.576 14.587 / 14.587
3D-Framework-185 14.595 / 14.584 14.585 / 14.582 14.572 / 14.595 14.584 / 14.587
3D-Framework-158 14.592 / 14.583 14.587 / 14.583 14.573 / 14.594 14.593 / 14.592
3D-Framework-518 14.583 / 14.585 14.592 / 14.584 14.573 / 14.576 14.594 / 14.594
3D-Framework-581 14.592 / 14.586 14.593 / 14.592 14.573 / 14.587 14.597 / 14.592
5.6 Evaluation of the Global explanation from the 3D-Framework and the XAI methods and the pattern learning results
To evaluate whether the global explanation from the 3D-Framework was superior to those provided by SHAP or GradCam, we scored the explanations with respect to faithfulness and complexity (see Table 1). For the simple-3D-MHL network, our proposed 3D framework outperformed total-GradCam and total-SHAP in terms of faithfulness score in the left hemisphere with values exceeding 0.21 compared to scores of less than 0.16 for total-GradCam, and less than 0.11 for total-SHAP. In the right hemisphere, our proposed 3D framework again outperformed total-GradCam and total-SHAP, achieving faithfulness scores over 0.18 compared to scores of less than 0.13 for total-GradCam, and less than 0.10 for total-SHAP. The 3D framework achieved the second-best result in complexity scores with total-GradCam having the lowest score in the left hemisphere and total-SHAP in the right hemisphere.
Up to this point, we have mainly explored the explanation results visually. However, we aimed to automate the process to identify the most significant subregions of interest based on the hypothesis (the classification task). To this end, we applied an affine registration to the total overlapping results of sulcal skeleton inputs from each hemisphere onto a probabilistic atlas of sulci (Perrot et al. (2011)). For this task, we explored the sulcal skeleton output of the simple-3D-MHL network as it slightly outperformed the simple-3D-CNN (see Supplementary Material Table 2 ii.) in the classification task and delivered better global explanations, faithfulness, and complexity scores than the simple-3D-CNN. Additionally, it follows patterns based on evidence from the literature (Fedeli et al. (2022); Harper et al. , Lopez-Persem et al. ).
Figure 7 presents the distribution of the most relevant voxels for outcome decisions within sulcal probabilistic areas according to different thresholds. The decisions in the right hemisphere were highly focused on specific sulci, with up to three sulci contributing to the 20% threshold, which we retained as the lower threshold (blue). Conversely, the left hemisphere predictions were based on broader considerations, with a number of sulci already contributing to the decision at the 5% threshold, which we retained as the lower threshold for the left hemisphere (blue).
For both conditions and both hemispheres, a specific focus was on the superior temporal sulcus and its posterior branches with a smaller but consistent contribution of the internal parietal and sub-parietal sulci. The noPCS condition on the left hemisphere additionally focused on the precentral sulcus and the Sylvian fissure. Interestingly, no specific focus was oriented towards the internal frontal sulcus (S.F.int), the probabilistic region in which the PCS is located when present.
a Refer to caption
b Refer to caption
Figure 7: A representation depicting the presence and absence of the paracingulate sulcus on the left hemisphere sulcal skeleton input a, is the histogram of the number of voxels per sulcus based on the probabilistic mapping of sulci for both conditions (PCS and noPCS) and both hemispheres, using sulcal skeleton brain input images with the simple-3D-MHL network. The voxels are extracted after thresholding for highest explainability values, with thresholds of the highest 5, 10, and 20% intensity on the left hemisphere (high threshold: blue, medium threshold: orange, low threshold: grey). b, the total overlapping pattern learning results of the three most significant sulcal sub-region for the three different level of intensity thresholding. The acronyms for all sulci are defined in Supplementary Extended DATA Fig. 5., and NA: undefined.
6 Discussion
Explainability is essential in medical imaging. Healthcare applications of AI need to able to explain their decision making to build trust and ensure that their predictions align with other symptoms and signs that affect health. Neuroimaging, the combination of brain images and computational methods, is a research application of medical imaging. Here, explainability for AI predictions support the assessment of the validity of results, but can also identify key contributors to decisions that in themselves reveal new patterns and directions for future investigation.
Our prior study (Mamalakis et al. (2024)) categorized the need for AI explanations into self-explainable, semi-explainable, non-explainable applications, and new-pattern discovery, based on the variability of expert opinions, the stability of the evaluation protocol and the representation dimensionality of the application. We applied the proposed guidance in a binary classification task related to a sub-region of the medial surface of the brain where secondary sulci are highly reproducible related with symptoms of psychosis, specifically hallucinations. This application was of the new-pattern discovery class. The output of the explainability indicated a wide distribution of brain regions on which the predictions depend suggesting covariant development of these regions during the perinatal period (De Vareilles et al. (2023)).
Automatic classification of psychotic and control patients based on structural MRI is a challenging task (Nunes et al. (2020)). In most cases, an acceptable detection rate is around 80.0% in clinical applications. However, in highly heterogeneous and variable cohorts, a lesser performance can be acceptable; approximately 60.0-70.0% acceptance accuracy; Iniesta et al. (2018); Hosmer and Lemeshow (2000). The TOP-OSLO cohort was particularly difficult for classification tasks of psychotic and control patients using structural MRI, with an accuracy of less than 60.0% (Nunes et al. (2020)). The variability of the paracingulate region in the TOP-OSLO cohort and the heterogeneity of the dataset create a highly challenging context for the automatic binary classification task of PCS presence. In this study, an accuracy of more than 70.0% in the left hemisphere and more than 60.0% in the right hemisphere was achieved, delivering an acceptable automated 3D deep learning network (Iniesta et al. (2018); Hosmer and Lemeshow (2000)) to apply global explainability methods for new-pattern discovery (Mamalakis et al. (2024)).
For the binary classification task of PCS presence, we developed two different 3D deep learning networks: a simple 3D convolutional neural network and a two-headed attention layer network. These networks utilized 3D brain inputs derived from preprocessed structural MRI scans, which included grey-white surface boundaries and sulcal skeletons from both hemispheres of a well-annotated cohort of 596 subjects. The performance of all networks was higher in the left hemisphere than in the right hemisphere. This discrepancy in performance was expected as the PCS is more prominent in the left hemisphere, including in psychopathological situations such as schizophrenia (Garrison et al. (2015)). Moreover, the left PCS has more associations with regional cortical thickness and sulcal depth than the right PCS, implying greater covariability of anatomical features in our input modalities with the presence of the PCS in the left hemisphere (Fornito et al. (2006)).
We developed an XAI 3D-Framework to fill the gap in the literature by providing accurate, low-complexity, three-dimensional (3D) global explanations in the neuroimaging field, where complex representational spaces demand more than basic two-dimensional interpretations. This framework was designed to offer robust and faithful global explanations for a deep learning network in the binary classification task of PCS presence. Our global explanations surpassed those produced by GradCam and SHAP in terms of faithfulness, providing a reliable interpretation of the deep networks for this classification task.
The overall explainability outputs cover wide regions of the brain, but we can notice some repetitive patterns through the different pipelines and modalities. In particular, there is a repeat of the cingulate region, the posterior temporal region, and both the medial and inferior frontal cortices. This may reflect some neurodevelopmental intertwining of the macroscopical development of the PCS and these regions. Regarding the cingulate and medial frontal regions, this intertwining is rather self-explanatory as the PCS is located in the medial frontal region, directly adjacent to the cingulate region, and as such the developmental events leading to the formation of a PCS are very likely to affect these regions. The two other regions mostly standing out are the inferior frontal region and the posterior temporal region. In the fetus, the sulci matching these regions (namely the inferior frontal sulcus and the posterior superior temporal sulcus) have both been reported to start appearing at 26 weeks of gestational age (w GA), while the cingulate sulcus appears earlier (around 23w GA) and the “secondary cingular sulci”, which encompass the PCS, appear later, at 31w GA (Garel et al. (2001)). This may point towards a time-window which is decisive to the development of the PCS, anterior to its actual apparition.
In terms of functional interpretation, it is interesting to notice a striking similarity between the regions on which the AI mostly focuses and the regions which have been reported to show the most functional connectivity with the paracingulate region (Fedeli et al. (2022); Harper et al. ), and the anterior cingulate region (Lopez-Persem et al. ). These studies report relevant functional connectivity between the medial frontal lobe (including the PCS) and the temporal region (including a focus on the posterior superior temporal region), the inferior frontal region, and the medial parietal region, which are all regions showing particular interest in the present work. Both Fedeli et al. (2022) and Harper et al. investigate the relation between the presence of a PCS and the related functional connectivity in these regions, and obtain more focused results (respectively in the cerebellum and superior anterior temporal region, or in the medial frontal region), but the important functional relationships between these regions support the relevance of the regions highlighted by our results.
The effectiveness of our framework remains dependent on the quality and diversity of accessible data, and its performance could differ in alternate contexts. Additionally, while the incorporation of explainability techniques contributes to interpretability, it might not encompass all underlying intricacies of network behavior. The selection of the two local methods necessitates further exploration of alternative approaches to ensure comprehensiveness. The enhancement of classification outcomes demands the incorporation of additional information beyond structural MRI data to achieve improved accuracy. To further investigate and incorporate more intricate strategies, the combination of XAI techniques and reduce dimensionality methods can explore overlapping aspects. Furthermore, complex approaches like t-SNE for non-linear dimensionality reduction can be employed to delve deeper into the subject.
Moving forward, we envision expanding the framework’s applicability to other neurological conditions and classification tasks (such as schizophrenia or bipolar disorder), employing other external datasets and further refining the interpretability techniques. Our study paves the way for a more systematic exploration of sulcal variability through the lens of deep learning, which can help explore cognitive and functional variability as well as pathological changes.
7 Conclusion
In this study, we classified the presence and absence of the secondary sulcus, the PCS, in a highly heterogeneous, unbiased, and well-annotated cohort of 596 structural MRIs. We used traditional 3D deep convolutional neural and attention mechanism networks to achieve acceptable detection accuracy and apply explainability techniques for new-pattern discovery.
As there were no global explainability methodologies in the literature efficient enough to deliver robust, accurate, low-complexity and three-dimensional explanations, we developed and proposed a novel XAI 3D-Framework to provide global explanations for the task of PCS detection. The global explanations derived from this framework outperformed established methods such as SHAP and GradCam in terms of faithfulness, achieving the quality needed to discover new global patterns. Additionally, the 3D representation of the proposed framework allows for the robust identification of complex new patterns within the high complexity and dimensionality of neuroscience applications. To accomplish this, the 3D-Framework enables the identification of sub-regions that are critical to decision-making through a post-fusion transformation of global explanations and statistical features.
The detection of the presence of the PCS appeared to focus most reliably on features related to the posterior temporal region and the internal parietal region (on the sulcal skeleton), with additional focus on the cingulate region and thalamus when focusing on the grey-white surface. This highlights potential co-variation between these structures which should be investigated further, notably in terms of common genetic influences in these regions, both for neurodevelopmental and functional or pathological considerations.
Significantly, our framework has revealed notable correlations between specific brain sub-regions and the presence or absence of the paracingulate sulcus, offering deeper insights into underlying developmental mechanisms. This approach opens the door to more robust and reliable explanations in medical imaging and other complex computer vision tasks.
This work advances the fields of both deep learning and neuroscience, enabling automated and impartial annotations while illuminating detailed neuroanatomical relationships across diverse subregions. Our XAI 3D-Framework offers unparalleled insights, thus setting the stage for comprehensive exploration within the ever-evolving domains of deep learning and neuroscience.
Acknowledgment
All research at the Department of Psychiatry in the University of Cambridge is supported by the NIHR Cambridge Biomedical Research Centre (NIHR203312) and the NIHR Applied Research Collaboration East of England. The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care. This study was supported by funding from the Medical Research Council, grant number: MR/W020025/1. We acknowledge the use of the facilities of the Research Computing Services (RCS) of University of Cambridge, UK. GKM consults for ieso digital health. All other authors declare that they have no competing interests.